### v3

# An Expanded Massive Multilingual Dataset for High-Performance Language Technologies (HPLT)

Abstract
Training state-of-the-art large language models requires vast amounts of clean and diverse textual data. However, building suitable multilingual datasets remains a challenge. In this work, we present HPLT v2, a collection of high-quality multilingual monolingual and parallel corpora, extending prior work of the HPLT project. The monolingual portion of the data contains 8T tokens covering 193 languages, while the parallel data contains 380M sentence pairs covering 51 languages. We document the entire data pipeline and release the code to reproduce it. We provide extensive analysis of the quality and characteristics of our data. Finally, we evaluate the performance of language models and machine translation systems trained on HPLT v2, demonstrating its value.
An Expanded Massive Multilingual Dataset for
High-Performance Language Technologies (HPLT)
Laurie Burchell1, Ona de Gibert2, Nikolay Arefyev3, Mikko Aulamo2, Marta Bañón4, Pinzhen Chen1, Mariia Fedorova3, Liane Guillou1, Barry Haddow1, Jan Hajič5, Jindřich Helcl5, Erik Henriksson6, Mateusz Klimaszewski1, Ville Komulainen6, Andrey Kutuzov3, Joona Kytöniemi6, Veronika Laippala6, Petter Mæhlum3, Bhavitvya Malik1, Farrokh Mehryary6, Vladislav Mikhailov3, Nikita Moghe7, Amanda Myntti6, Dayyán O’Brien1, Stephan Oepen3, Proyag Pal1, Jousia Piha6, Sampo Pyysalo6, Gema Ramírez-Sánchez4, David Samuel3, Pavel Stepachev1, Jörg Tiedemann2, Dušan Variš5, Tereza Vojtěchová5, Jaume Zaragoza-Bernabeu4 1University of Edinburgh, 2University of Helsinki, 3University of Oslo, 4Prompsit Language Engineering, 5Charles University, 6University of Turku, 7Amazonthanks: Work was done prior to joining Amazon. Contact: https://hplt-project.org
1 Introduction
In order to train the state-of-the-art large language models (LLMs) required for modern NLP, large amounts of high-quality textual training data are essential. However, obtaining a sufficient quantity of such data is far from easy. In addition, effective NLP research requires open training data so that results can be replicated and verified.
In this paper, we introduce a new set of text corpora dubbed HPLT v2.111https://hplt-project.org/datasets/v2.0 It is extracted from 4.5 petabytes (PB) of the Internet Archive (IA) [^2] https://archive.org and Common Crawl (CC) [^3] https://commoncrawl.org data. We build on the work of de Gibert et al. (2024) (hereafter referred to as HPLT v1.2) with an improved extraction pipeline and a much larger set of input crawls to produce the HPLT v2 collection of monolingual and parallel corpora. To our knowledge, our new corpus is the only large-scale text collection extracted from the IA, apart from HPLT v1.2. We release HPLT v2 under the permissive Creative Commons Zero (CC0) license [^4] https://creativecommons.org/share-your-work/public-domain/cc0/. We do not claim ownership of any of the text from which this data has been extracted. and provide the code to replicate our pipeline. Our main contributions can be summarised as:
  • We release monolingual corpora covering 193 languages and containing approximately 52 trillion characters and 8 trillion tokens.
  • We derive parallel corpora from our monolingual data for 50 languages paired with English, containing over 380 million sentence pairs.
  • We make the tools and pipelines used to create the collection openly available. [^5] https://github.com/hplt-project/HPLT-textpipes
  • We conduct an in-depth analysis of our data including descriptive statistics, manual inspection, and automatic register labelling.
  • We demonstrate the quality of HPLT v2 by using it to train a range of high-performing language and machine translation models.
Figure 1: The distribution of documents in the HPLT v2 cleaned dataset by language family and variety. Shortened ISO 639-3 language codes are used here; see the full plots at https://hplt-project.org/datasets/v2.0
2 Related work
The increasing data demands of state-of-the-art large language models have driven a rapid growth in both the number and the size of text corpora. We provide a summary of some well-known collections in Appendix A. Whilst LLMs trained on ostensibly English data have shown impressive multilingual capabilities Armengol-Estapé et al. (2022), of particular relevance to this work is the growing shift towards explicitly multilingual corpora. Compared with earlier efforts (e.g. OSCAR Suárez et al. (2019), CC-100 Conneau et al. (2020) and mC4 Xue et al. (2021)), more recent multilingual datasets cover increasing numbers of languages, e.g. CulturaX Nguyen et al. (2024a) and MADLAD-400 Kudugunta et al. (2024). HPLT v2 continues this trend by aiming for significant coverage of a wide range of languages. We note that the majority of previous multilingual datasets are sourced from CC, whereas much of HPLT v2 is composed of IA crawls. This means that HPLT v2 can be used in conjunction with these existing datasets as a complementary source.
Producing large-scale datasets by crawling the Web is helpful for scale, but raises questions around dataset quality such as the prevalence of boilerplate, explicit material or non-linguistic content Kreutzer et al. (2022). One way to tackle low-quality data is through human audit and curation (e.g. ROOTS Laurençon et al. (2022), Glot500-c Imani et al. (2023), Serengeti Adebara et al. (2023) and the MaLA Corpus Ji et al. (2024)). However, such an approach is difficult to scale. Instead, we ensure the quality of HPLT v2 through a robust dataset construction pipeline (Section 4) and by verifying our data through extensive analysis and downstream evaluation (Sections 5 and 6).
In addition to large-scale monolingual data in multiple languages, HPLT v2 contains high-quality parallel data. Whilst causal language models with decoder-only architectures rely primarily on monolingual data, recent studies have shown that incorporating parallel data during the pretraining stage significantly boosts multilingual, cross-lingual and machine translation (MT) performance for such models Kale et al. (2021); Briakou et al. (2023); Alves et al. (2024). Because of this, we expect that there is still significant demand for parallel data.
This work is a direct successor of our HPLT v1.2 dataset introduced in de Gibert et al. (2024). Compared to HPLT v1.2, we feature more data (21 billion vs. 5 billion documents) using an improved pipeline (Section 4), resulting in a significantly larger dataset (52 trillion characters compared to 42 trillion). The HPLT v2 collections are of higher quality than those in HPLT v1.2, as shown through comparative analysis and evaluation (Sections 5 and 6). For clarity, we list the main differences between HPLT v1.2 and HPLT v2 here:
  • The size of the source web collections is 2.5x larger, totalling 4.5 petabytes of compressed web data.
  • The text extraction pipeline uses Trafilatura rather than warc2text, which results in more efficient boilerplate removal.
  • Language identification uses a modified version of OpenLID rather than CLD2, increasing coverage from 75 to 193 language varieties.
  • Documents are annotated with their compliance to the robots.txt files of the corresponding websites, which can be used to filter out documents explicitly forbidden for crawling by website owners. The cleaned variant of the HPLT v2 dataset contains only robots.txt compliant documents.
  • Deduplication is done at collection level rather than globally.
  • Documents are annotated for PII information.
  • Document quality scores computed with web-docs-scorer rather than segment-level language model based scores
  • Simpler and more interpretable filtering and cleaning criteria.
3 Dataset description
In this section, we describe the HPLT v2 collection of monolingual and parallel corpora, before explaining how it was constructed in Section 4.
3.1 Monolingual datasets
The monolingual portion of HPLT v2 covers 193 language varieties [^6] Language varieties are labelled with an ISO 639-3 code denoting the variety plus an ISO 15924 four-letter code denoting the script, separated by an underscore: e.g., gla_Latn. and is published in two variants: ‘deduplicated’ (21 terabytes) and ‘cleaned’ (15 terabytes). In the latter variant, the documents filtered by our cleaning heuristics (see Section 4.2) are excluded. For training LLMs, we recommend using the cleaned variant, but we also publish the datasets before cleaning (‘deduplicated’) so that it is possible to apply custom cleaning pipelines to the HPLT v2 data. In total, the deduplicated monolingual HPLT v2 datasets contain approximately 7.6 trillion white-space separated tokens and 52 trillion characters, extracted from 21 billion documents. HPLT v2 is published in the JSONL format, with one document per line.
Figure 1 shows the distribution of documents in the cleaned monolingual data by language families and language variety. Indo-European languages, and especially English, make up the majority of the data. Unfortunately, this is the reality of current web crawls; increasing the amount of data available for other languages is not an easy task and is important future work. Table 6 gives a full breakdown of the statistics of the monolingual data.
Raw Filtered
Pairs Eng. words Pairs Eng. words
Total 1277M 16849M 380M 6780M
Median 11M 170M 4M 80M
Table 1: Counts in millions (M) of sentence pairs and English words in the parallel HPLT v2 data before filtering (Raw) and after filtering and deduplication (Filtered), both in total and the median across all languages.
3.2 Parallel datasets
We extract parallel data from the monolingual HPLT v2 to cover 50 languages paired with English. We aimed for a diverse range of language varieties and scripts in the low to medium resource range (listed in Table 5). We align these to English since this configuration has the highest potential for finding high-quality parallel data. We release our data in both XML and bitext format. [^7] https://opus.nlpl.eu/HPLT/corpus/version/HPLT
Table 1 gives the number of sentence pairs and English words per language prior to filtering (Raw) and after processing (Filtered). We provide both the total over the entire dataset and the median count by language variety. Our results show that the deduplicated HPLT v2 parallel corpora have a 70% reduction in sentence pairs compared to the raw data. The final dataset contains over 380 million sentence pairs, with the English side of the dataset containing over 6 billion words. The median number of sentence pairs by language variety is 4 million, but individual sizes vary greatly by language: the smallest, Sinhala, contains around 273 thousand pairs, whereas the largest, Finnish, contains over 29 million pairs. We give full statistics for each included language variety in Table 5 in the appendix.
We assume the large number of Finnish sentence pairs is due to the pipeline’s bias toward European languages. In contrast, languages such as Japanese and Korean, which we would expect to have larger corpora, may have lower counts because of lower-quality monolingual data and limited support in key pipeline components such as sentence splitting and tokenization. This results in reduced yields during data cleaning and filtering for non-European languages written in non-Latin scripts.
DocHPLT v2
We also introduce context, by providing the documents that our bitext data comes from. These documents are annotated with both sentence and paragraph alignment. Overall, we provide 74,078,581 aligned documents covering all 51 languages. This ranges from 38,199,407 documents in English to 20,448 documents in Xhosa. [^8] https://opus.nlpl.eu/legacy/DocHPLT.php
MultiHPLT v2
We leverage the English-centric HPLT v2 parallel resources to further create a multi-way parallel corpus, obtained by pivoting via English sentences. This corpus includes 1275 language pairs and contains over 16.7 billion parallel sentences. [^9] https://opus.nlpl.eu/MultiHPLT/corpus/version/MultiHPLT
A similar pivoting is also done for DocHPLT and we create the additional bitexts through sentences aligned to the same English document. However, anchoring sentences in aligned documents creates a different subset across the pivoted language pairs.
4 Dataset construction
In the following section, we explain the dataset construction pipeline for HPLT v2. We first extract text from web crawls via HTML (Section 4.1), deduplicate and clean this monolingual text (Section 4.2), and finally extract and process the parallel data (Section 4.3). Figure 2 provides a high-level overview of the pipeline.
4.1 Text extraction from web crawls
Sources
In total, we ingest 4.5 PB of web crawl data to build HPLT v2. 3.7 PB is sourced from IA from crawls conducted mostly between 2012 and 2020, with the remaining 0.8 PB coming from CC. We use CC crawls conducted mostly between 2014 and 2022. A detailed description of the crawls we use is in Appendix C.
Extracting HTML
Both IA and CC crawls are provided in the Web ARChive (WARC) format [^10] https://www.iso.org/standard/68004.html which stores HTTP requests and responses between a web crawler and web servers. We use the warc2text tool [^11] https://github.com/bitextor/warc2text to extract HTML and related metadata from these WARC files. It selects relevant WARC records containing HTML pages, removes documents from a list of known trash websites, [^12] Mostly containing auto-generated lists of phone numbers, addresses, etc.: https://github.com/paracrawl/cirrus-scripts/blob/master/url-filter-list.annotated and finally saves the results in the ZSTD-compressed JSONL format. The extracted metadata includes document URLs, paths to the original WARC files and record positions inside, content types and timestamps. Additionally, WARC records with URLs ending with "robots.txt" are stored for later use in filtering.
Extracting text
This stage of the pipeline extracts the main textual content from HTML pages and groups it into language-specific subsets. It first parses the HTML pages into a tree representation. Next, it removes likely machine-translated texts by searching for indicative HTML tags and attributes. It then removes boilerplate (i.e. parts of a web page that do not contribute to its main content) using Trafilatura 1.8.0 (Barbaresi, 2021). Following hyperparameter experimentation, we set include_comments=False, include_tables=False, no_fallback=False and MIN_EXTRACTED_SIZE=0, with all other hyperparameters set to their defaults. We chose not to use fallback to Trafilatura’s simple extraction baseline since it leaves most boilerplate intact, and we preferred sacrificing some documents but avoiding extra boilerplate in HPLT v2. Finally, we predict the language of the text using a modified version of the OpenLID model [^13] https://data.statmt.org/lid/lid193_merged_arabics.bin (Burchell et al., 2023; Burchell, 2024), where the Arabic dialects are combined under one macrolanguage label and the model training data has undergone improved pre-processing. These changes are intended to improve classification reliability. After text extraction, the dataset size reduces to 62 TB, 15 times smaller than the HTML data and 75 times smaller than the original web crawls.
Figure 2: Overview of the data acquisition and processing pipeline for HPLT v2.
4.2 Monolingual text processing
Following text extraction, we proceed to monolingual text cleaning in which we apply various criteria to select the cleanest documents. With the exception of fixing encoding, we do not alter the text in this process.
We first discard all documents for which the predicted probability of the language label is $<0.5$. We then perform crawl-level deduplication with a MinHash index Broder et al. (1998), using 240 hashes and a Jaccard similarity threshold of 0.8. We keep one document from each computed disjoint-set Galler and Fischer (1964), thus removing near-duplicates within each crawl.
To respect robots.txt [^14] https://www.robotstxt.org/ rules specified by each domain, we use the extracted robots.txt records to identify patterns disallowing the crawlers we use.151515*, CCBot, ia-archiver We use the fst [^16] https://burntsushi.net/transducers/ tool to create a compressed index of URLs to exclude and use it to remove documents originating from these URLs.
We then use a range of heuristics to discard low-quality documents. We calculate a document quality score using Web Docs Scorer (WDS), [^17] https://github.com/pablop16n/web-docs-scorer/ discarding documents with a score $<5$. We remove any documents where the length of the document is $<500$ characters, or where the average number of words per segment is $<5$ ($<10$ characters for Japanese, Chinese or Korean). We also filter documents where the URL is in the UT1 adult list. [^18] https://dsi.ut-capitole.fr/blacklists
Finally, we enrich documents with additional metadata. We add a unique identifier for the document hash derived from the WARC file name, the URL and the timestamp. We also carry out segment-level language identification (LID) using the Rust port [^19] https://github.com/ZJaume/heliport of HeLI-OTS Jauhiainen et al. (2022), trained on the OpenLID dataset. Finally, we add the Unicode character offsets of any personally identifiable information found by the PII tool. [^20] https://github.com/mmanteli/multilingual-PII-tool
Although the CC crawls are less than 20% of the input data, they are the source of about 60% of the final text. This is likely because CC focuses on textual content whereas IA includes much multimedia content, resulting in 4-8x lower yields in general. However, for some languages (e.g. Chinese, Persian, and a few smaller languages), IA provides more texts than CC. Appendix D presents a detailed study of the contributions of different source crawls to the final dataset.
4.3 Parallel data extraction
Our parallel data extraction pipeline is adapted from Bitextor. [^21] https://github.com/bitextor/bitextor We make the following changes to increase the quality of the final dataset:
  • Input data comes from cleaned monolingual HPLT v2 rather than WARCs.
  • We use Loomchild, a SRX-based sentence splitter Miłkowski and Lipski (2011), to cover more languages.
  • During sentence splitting, paragraph and sentence identifiers are added as persistent metadata through the pipeline.
  • Minimal length rule and fluency filtering in bicleaner-hardrules are disabled as this duplicates other processing steps.
  • Bicleaner AI (Zaragoza-Bernabeu et al., 2022) uses a multilingual model able to handle unseen language pairs during training.
  • Document-level output from the document-matching step is collected to allow the creation of document-level parallel data.
To avoid the possible introduction of new bugs in the pipeline, given that many of the steps in it are made for 2-letter language codes, we convert 3-letter language codes to 2-letter before processing.
5 Data analysis
In this section, we present our analysis of the HPLT v2 data based on indirect quality indicators, manual inspection, and register labels.
5.1 Indirect quality indicators
We consider two types of indirect quality indicators: descriptive statistics and website domains.
Descriptive statistics
We calculate descriptive statistics for HPLT v2 using the HPLT Analytics tool. [^22] https://github.com/hplt-project/data-analytics-tool We compare to the cleaned HPLT v1.2 dataset (our previous release).
For the monolingual data, there are far more unique segments (22.2% of HPLT v1.2 vs. 40.9% of HPLT v2) but far fewer documents longer than 25 segments (90.8% vs. 23.2%). Similarly, the proportion of short segments is reduced (39.6% vs. 13.3%). These changes can be attributed to the use of Trafilatura and WDS. More segments match the document language (58.6% vs. 81.5%), driven by improvements in LID accuracy and a more aggressive document filtering strategy. Finally, we identify frequent $n$-grams and find that a substantial amount of textual boilerplate remains, particularly from Wikipedia and blogging platforms.
Regarding parallel data, we find that the number of source and target tokens per language pair is much higher in HPLT v2 (by 47% and 49% on average) than in HPLT v1.2. Furthermore, 80% of the sentence pairs have a translation likelihood score from 0.8 to 1 (as computed by Bicleaner AI) which attests to their high quality. The frequent $n$-grams in the parallel datasets are similar among all languages: larger datasets tend to focus on hotels and legal notices, whereas the smaller datasets exhibit more variety and the frequent $n$-grams in these datasets reflect local content likely from news websites, e.g. political figures and place names. Appendix D contains further examples.
Domains
We explore the website domain names and geographic top-level domains present in the data in order to understand its origins better.
We find different patterns of website domain names in the corpora depending on the size of the language dataset. Languages with more data available contain a diverse range of website domain names in the monolingual data but more travel-related webpages in the parallel data. However, smaller language datasets tend to contain more Wikipedia and religious content in both the monolingual and parallel data. Appendix F contains further information about common domains.
Whilst most of the TLDs in our dataset are general purpose (e.g. .com, .org), we found that the most common geographic TLDs in the monolingual language corpora were usually from the country with the most speakers. This gave us confidence in the reliability of the text. We found the proportion of geographic TLDs from an indicative country was highest for those in Europe, whereas the datasets for many languages primarily spoken in Africa mostly consisted of general-purpose TLDs. The parallel data exhibits more diversity in TLDs than the monolingual data. For example, .eu is much more frequent, appearing in the top-10 TLDs of all mid-size and large parallel datasets of nearly all European languages. A more detailed discussion of our observations is in Appendix G.
5.2 Manual data inspection
To assess human-perceived quality, we manually inspected a random sample of documents from the cleaned monolingual datasets in 22 languages. Specifically, for each language spoken by the authors, we sampled 50 random documents extracted from each of the four groups of crawls: the older CC/IA crawls from 2012–2014, and the newer CC/IA crawls from 2017–2020. The main goal of this stratification was to compare the quality of texts we get depending on the crawl source and age, and select the most promising crawls for the next release of our datasets.
We asked participants to annotate any documents which look like pornographic content, look unnatural, and/or are not in the target language. Appendix H describes the inspection procedure and the results. Overall, for most languages, both the proportions of pornographic content and texts not in the target language are around 0–3%, with no significant difference between groups of crawls. Asturian, Scottish Gaelic and Norwegian Nynorsk are notable exceptions, with 31, 11 and 7 percent of texts not in the target language respectively. The proportion of unnatural texts is around 10% on average and up to 30% for some languages, leaving space for improvements. We also observe that the probability of getting an unnatural text from the newer CC crawls is roughly half of that of the other three inspected groups of crawls. This is probably related to the introduction of harmonic centrality ranking for domain prioritization in the CC crawler queue since 2017 (Nagel, 2023), which is stated to be more efficient in avoiding spam compared to the previously used techniques.
5.3 Register labels
As noted in Section 5.1, web crawls cover a vast range of different kinds of documents from various sources. We use automatic register (or genre) classification to create metadata about this variation, allowing users to make informed decisions when sampling from the data.
We use the multilingual register classifier described in Henriksson et al. (2024) to label the entire monolingual HPLT v2. This classifier covers 16 languages and is based on an XLM-R Large model (Conneau et al., 2020), fine-tuned on a multilingual web corpus manually annotated with register information. The classifier employs a hierarchical taxonomy with 25 register classes organized into 9 main categories (listed in Table 2). This label scheme is specifically designed for the linguistic characteristics of web texts and is therefore well-suited for our dataset.
The system achieves a mean micro F1 score of 77% on the 5 languages used during fine-tuning. It also demonstrates good performance for 11 unseen languages, with a mean micro F1 score of 66%. These results allow us to extend register labelling to a broad range of languages, though we limit predictions to languages within the 100 languages covered by XLM-RoBERTa.
We provide the classification certainty as well as the label, so that the threshold can be optimized by use case. Table 2 presents the distribution for register labels in our English data for a classification threshold of 0.4. Further work could use our derived labels to improve dataset quality, by e.g. filtering out MT content.
Register Percentage
How-to Interactive (HI) 1.8 %
Interactive Discussion (ID) 6.5 %
Informative Description (IN) 27.1 %
Informative Persuasion (IP) 10.8 %
Lyrical (LY) 0.5 %
Machine Translated (MT) 3.3 %
Narrative (NA) 18.1 %
Opinion (OP) 5.4 %
Spoken (SP) 0.2 %
Multiple labels 23.6 %
No label 2.5 %
Table 2: Register label distribution in HPLT v2 English dataset for classification threshold 0.4. See Henriksson et al. (2024) for the full scheme and explanation of the contents of the classes.
6 Empirical evaluation
In this section, we describe our empirical evaluation of the quality of the HPLT v2 monolingual corpora. We conduct this evaluation by employing the datasets as training material for several natural language processing (NLP) models.
6.1 Basic linguistic tasks and MLMs
We train masked language models on 52 different languages from the HPLT v2 datasets, choosing those with available benchmarks. [^23] https://github.com/hplt-project/HPLT-WP4 We use LTG-BERT (Samuel et al., 2023) to allow comparison with HPLT v1.2. We give full details about LTG-BERT in Appendix J.
We evaluate the trained MLMs on part-of-speech tagging, lemmatization and dependency parsing using the Universal Dependencies (UD) treebanks (de Marneffe et al., 2021), as well as named entity recognition (NER) using WikiAnn datasets (Pan et al., 2017). We compare to mBERT (Devlin et al., 2019) and XLM-R (Conneau et al., 2020) models as multilingual baselines, and to HPLT v1.2 BERT models [^24] https://hf.co/collections/HPLT/hplt-bert-models-6625a8f3e0f8ed1c9a4fa96d as monolingual baselines. The performance is measured using the official CoNLL 2018 evaluation code (Zeman et al., 2018) for the UD tasks, and seqeval Nakayama (2018) balanced F1 score for the NER task.
Figure 3: Win rates for MLMs at part-of-speech tagging, lemmatisation, dependency parsing, and named entity recognition.
Figure 3 shows the win rates achieved by the models for the four tasks (‘win rate’ here is the number of languages on which a given model outperforms other models). Models trained on the HPLT v2 datasets show a considerably higher win rate compared to the baselines in all the tasks except lemmatization, where XLM-R and HPLT v1.2 yield competitive results. However, we note that the difference between XLM-R, HPLT v1.2 and HPLT v2 on the lemmatization task is less than 1% of accuracy, meaning that no model significantly outperforms any other. Detailed scores by language and task are to be found in Table 13. We make the HPLT v2 BERT models with intermediate checkpoints publicly available. [^25] https://hf.co/collections/HPLT/hplt-20-bert-models-67ba52ae96b1fb8aae673493
6.2 NLU tasks and large generative LMs
Pretraining generative language models and evaluating their downstream performance on advanced natural language understanding (NLU) tasks is an established way to measure and compare training data quality Gao et al. (2020); Penedo et al. (2023); Longpre et al. (2024). Following Penedo et al. (2024a), we compare various large web-crawled pretraining corpora using this method for one high-resource and one low-resource language: English and Norwegian. We train 1.7B decoder-only LMs using 100B/30B tokens sampled from the English/Norwegian parts of our HPLT v2 dataset respectively. We compare our English and Norwegian models with models trained on same-sized samples of HPLT v1.2 de Gibert et al. (2024) and FineWeb Penedo et al. (2024a), and additionally compare our Norwegian models with FineWeb-2 Penedo et al. (2024b), CulturaX Nguyen et al. (2024b), and mC4 Xue et al. (2021). We replicate the design by Penedo et al. (2024a) and train the models with a fixed pretraining setup except for the pretraining corpus (English: four corpora; Norwegian: five corpora). We provide full details on pretraining and evaluation in Appendix I and describe our key results below.
Figure 4: Performance comparison of the trained generative LMs on English.
English
Average results over the English benchmarks are presented in Figure 4. Our models trained on the cleaned HPLT v2 datasets reach similar performance to the models trained on FineWeb data in downstream tasks, and considerably outperform the models trained on HPLT v1.2 and on the deduplicated subset of HPLT v2. This implies that our cleaning approach has successfully improved the data quality with respect to these benchmarks.
Norwegian
Average normalized scores over the Norwegian tasks are shown in Figure 5. We observe that the Norwegian models trained on FineWeb, CulturaX, and mC4 perform on par with HPLT v2 and outperform those trained on HPLT v1.2. Performance gains start to level off after 16B tokens, with the FineWeb and HPLT v2 scores being more stable during pretraining. This suggests that CulturaX, FineWeb, and HPLT v2 are more effective corpora for Norwegian, and their mixtures potentially provide further benefits.
Figure 5: Performance comparison of the trained generative LMs on Norwegian.
6.3 Machine translation tasks
We extrinsically evaluate the quality of the HPLT v2 parallel data by measuring the performance of MT models trained on it. This section considers two settings:
  • We compare HPLT v2 to its predecessor, HPLT v1.2, where we measure their performance in 10 overlapping languages.
  • We assess HPLT v2 as a complementary dataset to existing MT resources. We compare models in three data conditions: 1) solely on HPLT v2 parallel data; 2) on the data from Tatoeba Challenge (which includes most of the OPUS collection for each direction Tiedemann, 2012, 2020); and 3) on a combination of the two.
We carry out training for each individual language pair (to and from English) using the Transformer-base architecture (Vaswani et al., 2017) and the Marian NMT toolkit (Junczys-Dowmunt et al., 2018). Data processing and training are streamlined with OpusPocus [^26] https://github.com/hplt-project/OpusPocus following the configuration of Arefyev et al. (2024). We evaluate all models on the FLORES-200 benchmark (NLLB Team et al., 2024) using BLEU (Papineni et al., 2002), chrF++ (Popović, 2017), and COMET-22-DA (Rei et al., 2022). We use sacrebleu’s implementation of BLEU272727nrefs:1|case:mixed|eff:no|smooth:exp|version:2.5.1,
and where applicable, tok:ja-mecab, tok:ko-mecab, or tok:13a
and chrF++282828nrefs:1|case:mixed|eff:yes|nc:6|nw:0|space:no|version:2.5.1 with signatures footnoted (Post, 2018). Full results for each metric and each direction are given in Table 14 and Table 15 in Appendix K.
xx-en en-xx BLEU COMET BLEU COMET HPLT v1.2 28.5 0.7943 24.4 0.7623 HPLT v2 32.7 0.8343 27.9 0.8137 HPLT v2 28.7 0.8144 23.4 0.7941 OPUS 29.6 0.8142 23.4 0.8074 HPLT v2+OPUS 30.5 0.8237 24.2 0.8083
Table 3: Average scores for the HPLT v1.2 and v2 comparison (top) and HPLT v2 as a complimentary resource to OPUS (bottom). Only numbers that are available for all models in a comparison are averaged.
We report average BLEU and COMET scores in Table 3. It is worth noting that we only average numbers for translation directions that are covered by all models in a comparison. The top half features a clear advantage of HPLT v2 over HPLT v1.2, reflecting the effect of our improved data extraction pipeline discussed in Section 4. Results in the bottom of Table 3 shows that HPLT v2 and OPUS perform comparably. However, combining the two datasets leads to improvements in both BLEU and COMET. This indicates that HPLT v2 contains non-overlapping content relative to existing OPUS corpora, making it a valuable complementary resource for MT.
7 Conclusions and future work
This paper introduced the HPLT v2 dataset, a large-scale multilingual collection of openly-available monolingual and parallel web-crawled data. Building on our previous HPLT v1 effort, we focused on improving the quality of data available for a wider range of languages, and we make our data processing pipelines publicly available for easy reuse. We presented extensive data analysis as well as intrinsic and extrinsic evaluation, demonstrating the value of HPLT v2 for various NLP tasks.
Further work will focus on expanding language coverage and data quality, particularly for under-served languages. We also plan to conduct more experiments on the document-level aligned parallel corpus.
Limitations
Like many large-scale corpora, the majority of the data in HPLT v2 is in Indo-European languages, especially English, and the parallel data is English-centric. To an extent, this is a result of the dominance of these languages in the source web-crawl data. In addition, the evaluation in the paper only covers a subset of the languages in HPLT v2 due to a lack of resources for all languages present. We hope that the data we release in multiple under-served languages will be used to improve language technologies for more communities.
Whilst we focus on improving the HPLT v2 data processing pipeline, there are still residual errors in the final dataset in LID, boilerplate removal (particularly Wiki* boilerplate) and other cleaning steps. We make the code for our pipeline available to facilitate its evaluation and improvement. We note that there is only limited removal of machine-generated content in HPLT v2 (i.e., content generated by technologies like MT and LLMs), as detecting such content remains a difficult task (Yang et al., 2024).
It is possible that some of the test data we use for evaluation is contained within HPLT v2 (for example, the Wikipedia-based test set for named entity recognition). Nevertheless, we believe that the results reported in Section 6 are still indicative of the quality of HPLT v2, since the large-scale datasets we compare against are likely to have similar contamination issues.
During the evaluation, we discovered that the punctuation for Chinese languages (and probably Korean and Japanese) in HPLT v2 had been normalised incorrectly to its Latin equivalent, causing a drop in measured performance for languages in this script. We will fix this in the next iteration of the HPLT pipeline.
Ethical considerations
We source our data from web crawls and since Internet text is largely unregulated, our final dataset may contain harmful content or amplify existing biases, despite the extensive filtering applied to mitigate these issues. One notable bias is the over-representation of religious content in smaller language corpora, which could lead to models trained on this data being biased towards this particular domain.
Another pressing ethical consideration is the significant environmental impact of producing large-scale datasets. We mitigate this impact by making the data openly available in multiple formats, limiting the need to reproduce the processing pipeline.
We report the estimated CPU and GPU cost in hours for our work to allow for more informed decision-making in future research efforts:
  • WARC to HTML extraction: 250K CPU
  • HTML to text extraction: 1.7M CPU
  • Monolingual data processing: 600K CPU
  • Parallel data cleaning and deduplication: 1.8M CPU and 23K GPU
  • Register labels classification: 36.7K GPU
  • MLM experiments: 1.8K CPU and 4K GPUs
  • Generative LMs training and evaluation: 21.5K CPU and 43K GPU
  • MT models training and evaluation: 20K GPU
The total amount of hours spent would be roughly 4.4M CPU hours and 106K GPU hours. The most expensive task is the evaluation of our data through generative model training. We mitigate the environmental impact of our work by using one of the most eco-efficient HPC clusters in the world (LUMI) to carry out much of our computation. The LUMI supercomputer uses renewable, carbon-neutral energy. [^29] https://lumi-supercomputer.eu/sustainable-future/
Acknowledgements
We acknowledge Tsz Kin Lam, Otto Tarkka, and Maryam Teimouribadelehdareh for their annotation work; we thank Jelmer van der Linde for his help with Go.
This project has received funding from the European Union’s Horizon Europe research and innovation programme under Grant agreement No 101070350 and from UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee [grant number 10052546]. The contents of this publication are the sole responsibility of its authors and do not necessarily reflect the opinion of the European Union.
The authors wish to thank CSC - IT Center for Science, Finland for computational resources and support. The project has also been supported by the Czech MEYS project No. CZ.02.01.01/00/23_025/0008691 and Research Infrastructure project LM2023062.
Appendix A Comparison of multilingual collections
Dataset Size (TB) Tokens (T) Langs % English Source English only The Pile Gao et al. (2020) 0.8 0.39 1 100 various (c.f Section 2) C4.en Raffel et al. (2020); Dodge et al. (2021) 0.3 0.16 1 100 Common Crawl RefinedWeb Penedo et al. (2023) 2.8 0.6 1 100 Common Crawl Dolma-Web Soldaini et al. (2024) - 2.28 1 100 Common Crawl FineWeb Penedo et al. (2024a) - 15 1 100 Common Crawl Multilingual OSCAR-23.01 Suárez et al. (2019) - 1.1 153 48.43 Common Crawl CC-100 Conneau et al. (2020) 2.39 0.3 100 18.84 Common Crawl mC4 Xue et al. (2021) - 6.3 101 5.67 Common Crawl ROOTS Laurençon et al. (2022) 1.6 0.4 46 30.03 BigScience Catalogue Data, Common Crawl, OSCAR Glot500-c Imani et al. (2023) 0.6 - 511 *2.16 various (c.f. Appendix C) Serengeti Adebara et al. (2023) 0.042 - 517 - various (c.f. Appendix C) CulturaX Nguyen et al. (2024a) 27 6.3 167 45.13 OSCAR, mC4 MADLAD-400-clean Kudugunta et al. (2024) - 2.6 419 50 Common Crawl MaLA (Ji et al., 2024) - 0.074 939 4 various (c.f. Section 2.1.4) monoHPLT v1.2-dedup de Gibert et al. (2024) 11 5.6 75 41 Common Crawl, Internet Archive HPLT v2 (monolingual, deduplicated) 21 7.6 193 44 Common Crawl, Internet Archive
Table 4: Comparison of selected massively multilingual collections of monolingual data listed in chronological order. We report size, token counts, language coverage, and the proportion of English content. - indicates that data is not available. * indicates that the English percentage was computed over sentence counts, instead of token counts.
Appendix B Parallel and monolingual data statistics
Raw Filtered TMX
Language Sentence Pairs English Words Sentence Pairs English Words Sentence Pairs English Words
sin_Sinh 929,844 15,647,062 450,122 8,248,007 273,430 5,932,234
npi_Deva 1,058,740 18,514,145 523,022 10,176,931 317,120 7,145,363
xho_Latn 1,223,514 16,524,728 655,790 9,339,359 405,605 5,998,358
mal_Mlym 1,686,113 25,600,791 795,653 12,475,940 547,168 9,656,086
nno_Latn 2,358,129 34,771,540 1,175,108 21,352,540 563,791 10,548,302
mar_Deva 2,067,311 34,952,324 952,116 19,606,305 656,962 15,113,175
guj_Gujr 2,134,977 38,906,708 1,165,483 23,631,881 716,777 16,564,683
kan_Knda 2,354,299 37,451,816 1,238,033 21,344,021 720,157 13,965,655
tel_Telu 2,924,532 46,227,504 1,513,237 25,963,464 902,962 17,487,796
tam_Taml 3,859,610 55,779,718 1,759,372 28,369,233 1,111,471 20,718,487
uzn_Latn 2,791,412 37,715,209 1,571,871 25,823,124 1,159,869 19,667,785
urd_Arab 3,866,815 101,346,427 2,200,602 65,830,839 1,399,893 47,591,409
eus_Latn 5,907,808 79,485,282 2,526,198 38,107,950 1,491,873 24,303,464
epo_Latn 5,664,237 91,081,114 3,190,135 60,141,119 1,521,821 30,986,721
mlt_Latn 7,434,717 114,046,030 2,651,758 50,044,197 1,529,471 32,243,598
kaz_Cyrl 3,827,170 55,027,673 2,628,328 39,138,283 1,943,935 30,216,073
swh_Latn 10,125,330 145,685,653 3,680,151 68,766,541 1,985,899 39,952,916
ben_Beng 6,376,109 106,303,435 3,920,955 70,350,081 2,328,136 49,851,040
isl_Latn 11,929,153 146,981,787 6,624,589 91,089,371 2,694,541 47,440,271
gle_Latn 7,685,880 133,441,028 4,421,130 89,932,030 2,697,582 59,065,530
glg_Latn 8,680,808 145,602,784 5,166,276 99,562,132 2,783,727 58,437,672
bel_Cyrl 11,493,046 154,657,914 6,092,481 90,760,902 3,140,958 50,113,002
azj_Latn 8,506,772 118,079,751 4,765,278 72,026,597 3,188,231 51,425,346
pes_Arab 9,434,306 192,718,387 5,391,049 130,840,005 3,448,296 95,822,037
cym_Latn 9,390,284 156,087,956 6,348,606 125,442,540 3,867,402 82,244,645
afr_Latn 15,901,372 246,703,185 7,452,216 139,249,006 3,987,340 80,857,410
mkd_Cyrl 10,815,504 185,651,668 7,175,217 131,839,062 3,991,617 78,629,993
tha_Thai 13,818,095 102,830,296 7,551,187 52,171,172 4,088,354 34,155,503
als_Latn 11,171,352 208,460,688 6,943,910 145,918,660 4,166,536 94,263,904
bos_Latn 12,480,871 193,998,734 7,527,232 139,457,685 4,559,328 92,723,229
srp_Cyrl 17,605,882 244,921,478 9,618,806 153,171,621 5,291,686 90,518,351
zsm_Latn 47,173,963 558,911,698 22,298,471 301,446,773 8,432,285 147,009,838
heb_Hebr 34,004,891 431,453,938 21,600,460 279,563,405 8,686,089 162,768,846
est_Latn 29,934,421 362,843,780 16,629,846 223,025,816 8,797,574 133,824,400
hin_Deva 26,345,062 500,967,390 16,337,324 372,581,817 9,926,620 263,709,932
slv_Latn 30,956,083 449,919,060 18,290,300 301,699,622 10,336,528 188,709,019
lvs_Latn 39,599,210 476,030,125 24,504,355 316,955,851 11,294,618 183,588,490
lit_Latn 47,035,968 553,786,167 27,879,310 351,354,617 12,881,354 205,285,778
cat_Latn 40,922,098 671,563,410 26,451,844 521,397,424 13,080,859 292,854,267
hrv_Latn 45,617,022 627,929,707 27,783,979 420,953,899 14,263,908 250,103,294
ara_Arab 41,671,896 759,192,353 31,389,602 618,611,620 17,505,366 424,460,900
kor_Hang 76,980,595 865,790,290 46,010,282 530,544,997 18,393,859 294,720,264
jpn_Jpan 105,291,263 575,207,953 48,568,992 178,462,644 18,894,019 80,778,230
vie_Latn 47,831,389 1,077,677,445 35,072,681 831,280,934 19,231,770 502,683,444
slk_Latn 62,840,882 798,323,342 40,704,524 566,060,898 20,056,339 332,818,300
tur_Latn 84,823,944 1,054,338,198 46,493,019 656,237,839 21,616,652 402,325,110
bul_Cyrl 63,059,982 939,666,726 40,936,972 670,141,259 22,725,326 420,768,808
nob_Latn 71,277,875 969,526,777 45,204,041 700,446,147 22,912,722 395,447,564
ukr_Cyrl 68,111,464 862,726,167 46,141,511 637,243,802 25,125,019 400,949,773
fin_Latn 98,138,078 1,028,494,366 59,836,942 667,840,953 29,067,875 383,463,787
Table 5: Statistics for the parallel portion of HPLT v2 before filtering (Raw), after Bicleaner AI (Filtered) and after deduplication (TMX). Languages are in increasing order of deduplicated sentence pairs.
Language Segments Tokens Characters Documents
ace_Arab 1.170e+02 8.363e+03 4.973e+04 1.600e+01
ace_Latn 2.062e+05 8.196e+06 5.083e+07 1.293e+04
afr_Latn 3.774e+07 1.000e+09 5.947e+09 1.457e+06
als_Latn 9.510e+07 2.713e+09 1.610e+10 5.385e+06
amh_Ethi 7.006e+06 1.959e+08 1.031e+09 2.955e+05
ara_Arab 2.200e+09 4.814e+10 2.795e+11 8.267e+07
asm_Beng 2.677e+06 7.344e+07 4.757e+08 1.757e+05
ast_Latn 7.426e+06 1.950e+08 1.244e+09 2.732e+05
awa_Deva 1.315e+05 6.049e+06 2.877e+07 7.281e+03
ayr_Latn 1.885e+05 3.068e+06 2.508e+07 9.223e+03
azb_Arab 2.389e+06 3.958e+07 2.602e+08 6.611e+04
azj_Latn 1.266e+08 2.569e+09 1.962e+10 6.485e+06
bak_Cyrl 3.139e+06 7.533e+07 5.585e+08 1.708e+05
bam_Latn 9.172e+04 3.982e+06 2.074e+07 5.721e+03
ban_Latn 6.011e+05 1.134e+07 7.724e+07 1.070e+04
bel_Cyrl 4.884e+07 1.212e+09 8.540e+09 2.320e+06
bem_Latn 1.335e+05 4.523e+06 3.232e+07 6.136e+03
ben_Beng 1.760e+08 4.639e+09 3.016e+10 1.104e+07
bho_Deva 4.583e+05 1.347e+07 6.865e+07 2.864e+04
bjn_Arab 1.953e+04 5.482e+05 3.317e+06 1.112e+03
bjn_Latn 3.663e+05 8.048e+06 5.597e+07 1.876e+04
bod_Tibt 4.650e+05 5.781e+06 2.685e+08 2.744e+04
bos_Latn 2.682e+08 7.255e+09 4.607e+10 1.461e+07
bug_Latn 3.855e+04 2.705e+06 1.931e+07 2.023e+03
bul_Cyrl 6.814e+08 1.530e+10 9.693e+10 2.809e+07
cat_Latn 3.833e+08 1.002e+10 6.019e+10 1.855e+07
ceb_Latn 2.865e+06 8.589e+07 5.157e+08 1.388e+05
ces_Latn 1.927e+09 4.208e+10 2.739e+11 7.529e+07
cjk_Latn 3.670e+04 9.647e+05 7.432e+06 1.196e+03
ckb_Arab 5.226e+06 1.426e+08 9.128e+08 2.737e+05
crh_Latn 1.381e+06 3.676e+07 2.811e+08 1.227e+05
cym_Latn 1.557e+07 4.090e+08 2.402e+09 7.581e+05
dan_Latn 8.730e+08 2.120e+10 1.334e+11 3.384e+07
deu_Latn 1.113e+10 2.515e+11 1.782e+12 4.821e+08
dik_Latn 3.465e+04 2.295e+06 1.154e+07 2.325e+03
dyu_Latn 2.456e+04 1.194e+06 5.552e+06 1.390e+03
dzo_Tibt 3.997e+04 4.222e+05 7.375e+06 1.626e+03
ell_Grek 1.849e+09 4.270e+10 2.835e+11 7.033e+07
eng_Latn 1.165e+11 2.862e+12 1.708e+13 4.389e+09
epo_Latn 2.035e+07 4.716e+08 2.976e+09 8.189e+05
est_Latn 2.644e+08 4.742e+09 3.602e+10 8.449e+06
eus_Latn 3.762e+07 7.767e+08 6.052e+09 1.974e+06
ewe_Latn 1.434e+05 4.308e+06 2.132e+07 3.772e+03
fao_Latn 4.526e+06 9.345e+07 5.818e+08 2.399e+05
fij_Latn 1.789e+05 7.263e+06 3.769e+07 8.914e+03
fin_Latn 9.766e+08 1.845e+10 1.557e+11 3.482e+07
fon_Latn 1.476e+04 1.233e+06 5.335e+06 1.226e+03
fra_Latn 1.056e+10 2.370e+11 1.457e+12 4.018e+08
fur_Latn 7.300e+05 2.082e+07 1.147e+08 3.667e+04
fuv_Latn 1.340e+05 5.143e+06 2.990e+07 7.760e+03
gaz_Latn 9.736e+05 2.888e+07 2.192e+08 4.914e+04
gla_Latn 3.307e+06 8.066e+07 4.836e+08 1.374e+05
gle_Latn 1.099e+07 2.957e+08 1.749e+09 4.908e+05
glg_Latn 6.118e+07 1.639e+09 1.011e+10 3.020e+06
grn_Latn 1.713e+06 3.072e+07 2.186e+08 7.342e+04
guj_Gujr 2.064e+07 5.768e+08 3.386e+09 1.134e+06
hat_Latn 4.635e+06 1.223e+08 6.389e+08 2.127e+05
hau_Latn 5.688e+06 1.526e+08 8.535e+08 3.159e+05
heb_Hebr 4.666e+08 9.966e+09 5.682e+10 1.712e+07
hin_Deva 2.674e+08 8.637e+09 4.396e+10 1.365e+07
hne_Deva 5.500e+04 2.199e+06 1.059e+07 2.806e+03
hrv_Latn 2.971e+08 7.307e+09 4.800e+10 1.230e+07
hun_Latn 1.419e+09 3.052e+10 2.252e+11 5.187e+07
hye_Armn 6.524e+07 1.405e+09 1.072e+10 3.599e+06
ibo_Latn 1.411e+06 3.829e+07 2.052e+08 5.629e+04
ilo_Latn 1.120e+06 2.478e+07 1.568e+08 4.875e+04
ind_Latn 2.389e+09 5.462e+10 3.842e+11 9.814e+07
isl_Latn 6.964e+07 1.536e+09 9.593e+09 2.841e+06
ita_Latn 5.127e+09 1.274e+11 8.206e+11 2.218e+08
jav_Latn 6.431e+06 1.378e+08 9.375e+08 1.960e+05
jpn_Jpan 2.327e+10 4.236e+10 9.011e+11 4.177e+08
kab_Latn 3.452e+05 9.222e+06 5.419e+07 1.510e+04
kac_Latn 1.594e+05 5.955e+06 2.840e+07 7.587e+03
kam_Latn 1.426e+04 6.740e+05 4.645e+06 1.183e+03
kan_Knda 2.493e+07 5.329e+08 4.298e+09 1.336e+06
kas_Arab 2.711e+04 6.780e+05 3.468e+06 9.490e+02
kas_Deva 1.357e+03 3.194e+04 1.854e+05 1.060e+02
kat_Geor 6.372e+07 1.244e+09 1.016e+10 3.335e+06
kaz_Cyrl 8.101e+07 1.409e+09 1.113e+10 2.637e+06
kbp_Latn 4.679e+04 4.258e+06 2.090e+07 7.075e+03
kea_Latn 4.391e+04 1.143e+06 6.144e+06 1.962e+03
khk_Cyrl 5.347e+07 1.342e+09 9.327e+09 2.121e+06
khm_Khmr 9.864e+06 1.138e+08 2.122e+09 7.010e+05
kik_Latn 5.193e+04 1.428e+06 9.292e+06 3.995e+03
kin_Latn 1.917e+06 5.074e+07 3.671e+08 9.270e+04
kir_Cyrl 1.004e+07 2.467e+08 1.925e+09 6.761e+05
kmb_Latn 1.180e+04 3.831e+05 2.068e+06 5.310e+02
kmr_Latn 7.147e+06 1.959e+08 1.123e+09 3.643e+05
knc_Arab 1.083e+04 2.620e+05 1.302e+06 2.450e+02
knc_Latn 1.052e+04 2.409e+06 1.195e+07 2.472e+03
kon_Latn 4.748e+04 1.944e+06 1.127e+07 2.542e+03
kor_Hang 1.358e+09 1.970e+10 8.923e+10 3.887e+07
lao_Laoo 3.200e+05 5.178e+06 8.468e+07 2.950e+04
lij_Latn 1.577e+05 5.593e+06 3.146e+07 8.371e+03
lim_Latn 7.140e+06 1.806e+08 1.125e+09 3.679e+05
lin_Latn 2.003e+05 5.555e+06 3.292e+07 7.588e+03
lit_Latn 3.222e+08 6.676e+09 5.039e+10 1.334e+07
lmo_Latn 2.125e+06 5.964e+07 3.454e+08 1.462e+05
ltg_Latn 1.514e+05 3.790e+06 2.688e+07 9.209e+03
ltz_Latn 5.059e+06 1.072e+08 7.104e+08 2.469e+05
lua_Latn 3.869e+04 1.368e+06 9.005e+06 1.083e+03
lug_Latn 4.075e+05 9.176e+06 6.796e+07 2.128e+04
luo_Latn 8.412e+04 3.727e+06 2.033e+07 4.153e+03
lus_Latn 3.433e+06 1.252e+08 6.520e+08 1.604e+05
lvs_Latn 1.738e+08 3.461e+09 2.518e+10 6.772e+06
mag_Deva 1.929e+04 8.906e+05 4.283e+06 3.280e+02
mai_Deva 6.455e+05 1.779e+07 9.674e+07 2.498e+04
mal_Mlym 4.800e+07 9.737e+08 9.489e+09 3.105e+06
mar_Deva 3.632e+07 9.807e+08 6.622e+09 2.080e+06
min_Latn 6.008e+05 1.098e+07 7.477e+07 2.504e+04
mkd_Cyrl 5.701e+07 1.485e+09 9.440e+09 3.566e+06
mlt_Latn 8.675e+06 1.958e+08 1.442e+09 3.673e+05
mni_Beng 6.576e+04 1.627e+06 1.179e+07 2.934e+03
mos_Latn 1.910e+04 8.075e+05 3.864e+06 9.310e+02
mri_Latn 2.795e+06 8.676e+07 4.243e+08 1.083e+05
mya_Mymr 3.050e+07 4.532e+08 5.819e+09 1.368e+06
nld_Latn 3.075e+09 7.141e+10 4.511e+11 1.387e+08
nno_Latn 3.460e+07 8.603e+08 5.404e+09 1.423e+06
nob_Latn 6.760e+08 2.154e+10 1.332e+11 2.705e+07
npi_Deva 3.714e+07 1.128e+09 7.256e+09 2.778e+06
nso_Latn 1.433e+05 5.322e+06 2.749e+07 6.066e+03
nus_Latn 8.514e+03 3.932e+05 1.882e+06 2.720e+02
nya_Latn 1.344e+06 2.706e+07 2.029e+08 5.312e+04
oci_Latn 4.195e+06 1.027e+08 6.354e+08 1.899e+05
ory_Orya 3.596e+06 1.201e+08 7.815e+08 4.129e+05
pag_Latn 8.583e+04 5.657e+06 3.352e+07 6.900e+03
pan_Guru 1.174e+07 3.722e+08 1.902e+09 5.846e+05
pap_Latn 1.387e+06 4.671e+07 2.541e+08 8.981e+04
pbt_Arab 8.455e+06 2.794e+08 1.304e+09 4.665e+05
pes_Arab 3.963e+09 8.855e+10 4.551e+11 9.050e+07
plt_Latn 4.736e+06 1.171e+08 8.103e+08 2.078e+05
pol_Latn 4.461e+09 8.953e+10 6.316e+11 1.754e+08
por_Latn 6.125e+09 1.463e+11 8.965e+11 2.378e+08
prs_Arab 6.900e+07 1.844e+09 9.567e+09 2.839e+06
quy_Latn 4.943e+05 1.731e+07 1.434e+08 3.694e+04
ron_Latn 1.697e+09 4.005e+10 2.507e+11 6.588e+07
run_Latn 1.752e+06 4.444e+07 3.165e+08 1.373e+05
rus_Cyrl 2.629e+10 5.409e+11 3.908e+12 8.847e+08
sag_Latn 5.190e+04 3.612e+06 1.674e+07 3.161e+03
san_Deva 3.281e+06 4.380e+07 3.592e+08 5.491e+04
sat_Olck 4.580e+04 1.085e+06 6.266e+06 2.566e+03
scn_Latn 1.650e+06 4.239e+07 2.523e+08 8.197e+04
shn_Mymr 9.214e+04 1.648e+06 2.121e+07 6.003e+03
sin_Sinh 3.371e+07 7.956e+08 4.981e+09 1.153e+06
slk_Latn 4.943e+08 1.063e+10 7.037e+10 2.183e+07
slv_Latn 2.386e+08 5.435e+09 3.526e+10 1.028e+07
smo_Latn 1.012e+06 3.709e+07 1.861e+08 4.586e+04
sna_Latn 1.202e+06 2.392e+07 1.926e+08 6.108e+04
snd_Arab 2.826e+06 8.953e+07 4.286e+08 1.003e+05
som_Latn 1.638e+07 3.888e+08 2.565e+09 9.665e+05
sot_Latn 1.085e+06 3.100e+07 1.715e+08 4.392e+04
spa_Latn 1.212e+10 3.220e+11 1.954e+12 5.031e+08
srd_Latn 9.171e+05 2.389e+07 1.487e+08 5.382e+04
srp_Cyrl 9.381e+07 2.519e+09 1.616e+10 4.123e+06
ssw_Latn 6.213e+04 9.943e+05 8.821e+06 2.036e+03
sun_Latn 3.238e+06 6.963e+07 4.753e+08 1.148e+05
swe_Latn 1.755e+09 4.011e+10 2.511e+11 6.681e+07
swh_Latn 3.431e+07 7.177e+08 4.664e+09 1.374e+06
szl_Latn 6.366e+05 1.468e+07 1.038e+08 4.093e+04
tam_Taml 1.686e+08 2.981e+09 2.624e+10 6.106e+06
taq_Latn 1.388e+04 1.544e+06 8.845e+06 1.747e+03
tat_Cyrl 1.345e+07 2.967e+08 2.157e+09 6.307e+05
tel_Telu 3.919e+07 8.354e+08 6.505e+09 2.058e+06
tgk_Cyrl 2.485e+07 6.248e+08 4.590e+09 1.261e+06
tgl_Latn 5.288e+07 1.346e+09 8.131e+09 1.869e+06
tha_Thai 3.391e+08 3.506e+09 5.998e+10 1.770e+07
tir_Ethi 1.128e+06 3.672e+07 1.816e+08 6.469e+04
tpi_Latn 2.824e+05 1.251e+07 6.453e+07 1.398e+04
tsn_Latn 1.322e+05 5.273e+06 2.767e+07 6.050e+03
tso_Latn 2.212e+05 8.668e+06 4.929e+07 1.101e+04
tuk_Latn 3.355e+06 7.068e+07 5.700e+08 1.710e+05
tum_Latn 9.901e+04 2.876e+06 2.110e+07 4.384e+03
tur_Latn 2.575e+09 5.167e+10 3.896e+11 1.166e+08
twi_Latn 1.256e+05 4.696e+06 2.418e+07 5.860e+03
uig_Arab 8.982e+06 2.239e+08 1.747e+09 4.424e+05
ukr_Cyrl 1.169e+09 2.523e+10 1.829e+11 4.740e+07
umb_Latn 5.991e+04 2.431e+06 1.541e+07 2.471e+03
urd_Arab 5.063e+07 2.126e+09 1.001e+10 3.194e+06
uzn_Latn 1.480e+07 3.513e+08 2.846e+09 7.069e+05
vec_Latn 1.579e+06 3.526e+07 2.180e+08 8.480e+04
vie_Latn 3.020e+09 8.320e+10 3.795e+11 1.007e+08
war_Latn 2.009e+05 5.889e+06 3.557e+07 1.387e+04
wol_Latn 1.615e+05 5.463e+06 2.754e+07 5.679e+03
xho_Latn 1.821e+06 3.034e+07 2.587e+08 6.309e+04
ydd_Hebr 2.940e+06 7.753e+07 4.585e+08 1.283e+05
yor_Latn 1.469e+06 4.281e+07 2.178e+08 6.613e+04
yue_Hant 1.235e+06 3.268e+06 7.430e+07 6.129e+04
zho_Hans 4.245e+10 7.403e+10 2.352e+12 1.247e+09
zho_Hant 4.480e+09 9.510e+09 2.868e+11 1.571e+08
zsm_Latn 5.798e+08 1.148e+10 7.843e+10 1.842e+07
zul_Latn 2.710e+06 4.436e+07 3.808e+08 1.136e+05
Table 6: Counts of segments, tokens, characters and documents for each language in the monolingual HPLT v2 datasets. Tokens are words as defined by Unix wc.
Appendix C Sources of web crawls
Name Years Size (TB)
IA full crawls 2012–2020 3390
wide5 2012–2012 365
wide6 2012–2013 204
wide10 2014–2014 91
wide11 2014–2014 420
wide12 2015–2015 449
wide15 2016–2017 358
wide16 2017–2018 768
wide17 2018–2020 641
survey3 2015–2016 94
CC full crawls 2014–2022 743
CC-MAIN-2014-35 2014 43
CC-MAIN-2014-42 2014 54
CC-MAIN-2015-11 2015 29
CC-MAIN-2015-48 2015 30
CC-MAIN-2017-04 2017 54
CC-MAIN-2018-05 2018 75
CC-MAIN-2018-22 2018 52
CC-MAIN-2018-43 2018 59
CC-MAIN-2021-43 2021 86
CC-MAIN-2022-27 2022 85
CC-MAIN-2022-40 2022 83
CC-MAIN-2022-49 2022 93
Partial crawls 2013-2023 317
1% of WARCs from 81 CC crawls 2013-2023 46
7% of IA ArchiveBot 2013-2023 271
Table 7: List of web crawls used to construct HPLT v2. From IA, we use 8 Wide crawls, 1 Survey crawl containing main pages of websites and a random sample of 7% of items from IA ArchiveBot. From CC, we use 12 randomly-selected full crawls, plus a 1% sample of WARCs from each of the other 81 available crawls.
Appendix D Yields of different crawls
Figure 6: Proportions of data from different groups of crawls at various processing stages. Crawls were quantified in TB of compressed WARC files, while raw texts and deduplicated cleaned texts in characters.
Figure 7: Yields (in characters per 1 GB of raw compressed crawls) of different crawls at different stages.
To figure out how different web crawls contribute to our datasets and which crawls are the most promising sources of monolingual corpora in general, we compared crawls from two points of view: the amount of texts extracted from each crawl and the quality of these texts. In this section, we study crawls from the first point of view, while in H the results of manual quality inspection are presented.
To make a comparison, we group all crawls into groups according to their age and source. The oldest IA wide crawls from 2012-2014 (from wide5 up to wide11) are assigned to the group ia_o, the newest wide16, wide17 crawls from 2017-2020 to the group ia_n, and the wide12, wide15 crawls in the middle to the group ia_m. CC crawls are split by age following the same time periods, but additionally a group cc_r is introduced for the recent CC crawls from 2021-2023 (we don’t have IA wide crawls from this time period). Finally, the IA survey3 and ArchiveBot crawls form their own groups ia_survey and ia_archivebot. In total, we have 9 groups of crawls.
For different processing stages, Figure 6 visualizes how much data comes from different groups of crawls. While originally less than 20% of our crawls are CC crawls, they contribute about half of the raw text before duplication and more than 60% of the text after deduplication and cleaning. Especially high-yielding are the new and recent CC crawls, they are only 6% and 8% of all crawls in size but contribute 28% and 30% of text (both when counting in characters and in documents) to the cleaned version. On the other hand, the newest IA wide crawls are 32% of all crawls in size but contribute only about 11% of text.
Figure 7 suggests another point of view showing yields for different crawls, or more specifically, how much text (measured in the number of characters) is extracted from 1 GB of compressed WARC files for each crawl. Evidently, CC crawls have the highest yields, especially the newer ones. Compared to the newer CC crawls, for the older CC crawls more data is filtered during deduplication and cleaning, giving finally lower yields despite a bit higher yields of raw texts. IA wide crawls have 4-8x smaller yields than CC crawls. The survey IA crawl has a comparable yield to the wide crawls in the final dataset. Since they are publicly available, it probably makes sense to employ more of these crawls in the future. Finally, the ArchiveBot IA crawl has remarkably low yields.
Figure 8: Proportions of texts from different groups of crawls for the 15 languages with the largest contribution of IA crawls.
Despite having a lower contribution in general, for some languages, IA crawls supply the majority of texts. Figure 8 shows 15 languages with the highest proportion of texts from IA crawls. They include both high-resourced (Chinese, Western Persian) and low-resourced languages. Deduplication and cleaning significantly reduce the number of languages with high contribution of IA. For instance, before deduplication and cleaning there are 49 languages having more than 70% of texts (characters) coming from IA and only 6 such languages after.
Appendix E Frequent n-grams
We obtain frequent $n$-grams (up to order 5) in each dataset after tokenizing text and applying some restrictions:
  • $n$-grams must start and end in the same segment (i.e. no line breaks are allowed in the middle of a $n$-gram)
  • $n$-grams containing any punctuation are discarded
  • $n$-grams that start or end in stopwords are discarded
  • $n$-grams are calculated case-insensitive
  • all tokens in the $n$-gram must have at least one alphabetic character
Figures 9 and 10 show the 5 most frequent $n$-grams (orders 1 to 5) in HPLT v2. In the case of parallel datasets, $n$-grams are selected from the target (non-English) side of the segments. Translation to English is obtained with Google Translate. [^30] https://translate.google.com/
We find that most datasets (both monolingual and parallel) contain frequent $n$-grams that seem to be boilerplate, such as "edit source", "read more", "click button" or "view map". This kind of content usually comes from Wikipedia and Blogspot. In the monolingual datasets, there is a large amount of text that seems to come from headers or footers in news webpages, e.g. "latest news". Biblical $n$-grams (such as "god" or "jehovah") are also very frequent in some datasets, notably African languages, matching our observations about frequent domains (Appendix F). Some frequent $n$-grams suggest poor-quality content in some datasets, since they seem to be related to downloads webpages, online game platforms or betting sites.
For the parallel datasets, we observe that on the English side the frequent $n$-grams are very similar across all languages. For the languages with the most data, hotels and legal notices are the most common kind of $n$-grams. The smaller parallel datasets tend to exhibit more variety of $n$-grams and include $n$-grams alluding to political leaders or city names, which suggest more locally-generated content (probably from news sites). Finally, frequent $n$-grams in parallel datasets from Eastern European countries usually contain mentions to European institutions (such as the European Parliament or the European Commission). This matches our observations on TLDs in Appendix G.
Figure 9: Frequent n-grams in monolingual datasets.
Figure 10: Frequent n-grams in parallel datasets (non-English side).
Appendix F Frequent domains
Inspecting the most common domain names in the datasets is one way to understand the type of content we can find in it. Table 8 gives the datasets with the highest proportion of frequent domain classes, and Table 9 gives the datasets with the highest proportion of frequent domain classes for the parallel data. We make the following general observations:
  • Mid-to-large-sized datasets show a wider variety of domains with no clear majority source. However, in monolingual datasets, blogging platforms usually get a significant portion of the total (Table 8).
  • Wikipedia tends to be among the most frequent domains for both monolingual and parallel datasets. It is usually the most frequent domain for smaller language datasets (Table 8).
  • Hotel and travel webpages are much more frequent in the larger parallel datasets and very infrequent in the monolingual data (Table 9).
  • News and media outlets are also a frequent content source in monolingual datasets, with some news websites getting a significant percentage in different datasets: for example, regional websites from the Free Radio Europe [^31] https://www.rferl.org/navigation/allsites or Voice of America [^32] https://www.voanews.com/navigation/allsites networks (Table 8).
  • Religious and biblical content is also very frequent in the smaller monolingual and parallel datasets. This is specially notable in the case of African languages, which often get more than three quarters of their content from such sites (Table 8).
  • Software and online gaming websites are usually among the top-10 most frequent domains in almost all parallel datasets.
  • Chinese shopping websites are common in the larger parallel datasets of non-European languages (Vietnamese, Japanese, Korean, Arabic and Turkish).
  • No pornographic webpages appear in the top domains, implying our filter for such content worked as expected.
Dataset % of documents Blogging platforms Standard Malay 50% Magahi 48% Greek 24% Cantonese 22% Portuguese 16% Finnish 16% Swedish 13% Spanish 10% Wikipedia Santali 90% Ligurian 80% Waray 74% Iloko 66% Esperanto 66% Occitan 62% Sicilian 55% News & media Crimean Tatar 61% Tigrinya 48% Banjar (Arabic) 46% Nigerian Fulfulde 38% Turkmen 32% Kyrgyz 30% Rundi 29% Religious & biblical Dyula 99% Fon 96% Bemba 95% Tumbuka 94% Kamba 93% Chokwe 92% Central Kanuri (Latin) 91% Luba-Lulua 88% Sango 88% Umbundu 84%
Table 8: Languages with the biggest proportion of frequent domain classes in the monolingual HPLT v2 corpora.
Language % of segments Hotels & travels Icelandic 42% Malay 34% Hebrew 26% Lithuanian 21% Korean 18% Thai 16% Norwegian Bokmål 16% Japanese 15% Wikipedia Norwegian Nynorsk 72% Galician 37% Esperanto 36% Kannada 22% Macedonian 21% Telugu 19% Catalan 19% Religious & biblical Xhosa 70% Esperanto 28% Swahili 20% Nepali 16% Icelandic 14% Albanian 14%
Table 9: Frequent domain classes in parallel HPLT v2 datasets for different languages (non-English side).
Appendix G Geographic TLDs
Tables 10 and 11 list the most frequent examples of geographic TLDs in the monolingual and parallel HPLT v2 corpora respectively. We make the following observations in addition to those made in the main text:
  • In general, the most frequent TLDs in many of the datasets are generic (such as .com, .org or .info).
  • Some TLDs are frequent because they "sound good" rather than indicating the kind of content or language: .icu (because it reads like "I see you"), .is (official TLD for Iceland, but used as a verb and very noticeably in bible.is, a religious webpage whose domain is usually in the top-10 most frequent TLDs), .tv (for the country of Tuvalu, but widely used for TV-related web pages), .co (for Colombia, but mostly used for companies), .no (for Norway, but used as a negative particle), .nu (for the island of Niue, but used because it sounds like "new"), etc.
  • There are common TLDs for super-national territories: .eu (European Union), .africa, .asia, etc.
  • In our monolingual datasets, there is frequently one geographic TLD among the 10 most frequent ones that clearly surpasses the others. The "winning" TLD is usually from the country where the dataset language is spoken most, indicating that the text content is probably in the correct language. The percentage of this "winning country" varies depending on the amount of general purpose TLD in the dataset, but it is in general higher for European countries. This "winning" geographic TLD, in the case of parallel datasets, is less frequent and, when present, its portion of the total is noticeably lower than for the monolingual datasets.
  • Many African languages do not have a significant portion of geographic TLDs (beyond the aforementioned .is, .no, etc).
  • For some languages, there are a few countries TLDs in the top-10 from closely related countries or territories (for example, from former colonial rulers (i.e. African languages datasets) or with geostrategic interests (i.e. .ru (Russia) appearing in all former Soviet states). This may indicate "language contamination" in the data.
Language % of TLD Country or documents Territory One geographic TLD Manipuri 80% .in India Lithuanian 79% .lt Lithuania Polish 77% .pl Poland Hungarian 76% .hu Hungary Danish 76% .dk Denmark Icelandic 73% .is Iceland Faroese 73% .fo Faroe Islands Macedonian 73% .mk N. Macedonia Latgalian 73% .lv Latvia Latvian 72% .lv Latvia Related territories Slovak 77% .sk Slovakia 3% .cz Czechia Kazakh 71% .kz Kazakhstan 3% .ru Russia Russian 65% .ru Russia 5% .ua Ukraine 2% .by Belarus Croatian 47% .hr Croatia 3% .ba Bosnia 3% .rs Serbia Kyrgyz 33% .kg Kyrgyzstan 7% .ru Russia Bosnian 29% .rs Serbia 13% .ba Bosnia Language variants Romanian 72% .ro Romania 3% .md Moldova Dutch 66% .nl Netherlands 12% .be Belgium German 60% .de Germany 6% .at Austria 5% .ch Switzerland Portuguese 45% .br Brazil 9% .pt Portugal Lombard 47% .ch Switzerland 5% .it Italy Uyghur 36% .cn China 3% .kz Kazakhstan French 30% .fr France 3% .be Belgium 2% .ca Canada 2% .ch Switzerland Spanish 15% .es Spain 4% .ar Argentina 4% .mx Mexico 2% .cl Chile 1% .pe Peru
Table 10: Frequent geographic TLDs in monolingual HPLT v2 datasets for different languages.
Language % of TLD Country or segments Territory One geographic TLD Norwegian Nynorsk 36% .no Norway Norwegian Bokmål 34% .no Norway Azerbaijani 33% .az Azerbaijan Macedonian 25% .mk North Macedonia Vietnamese 23% .vn Vietnam Farsi 22% .ir Iran Hebrew 20% .il Israel Sinhala 19% .lk Sri Lanka Serbian 16% .rs Serbia Malay 15% .my Malaysia Hindi 15% .in India Japanese 15% .jp Japan Korean 15% .kr South Korea Related territories (European Union) Maltese 68% .eu European Union 4% .mt Malta Slovene 31% .si Slovenia 17% .eu European Union Estonian 35% .ee Estonia 16% .eu European Union Latvian 30% .lv Latvia 16% .eu European Union Lithuanian 35% .lt Lithuania 14% .eu European Union Slovak 44% .sk Slovakia 11% .eu European Union 4% .cz Czechia Croatian 26% .hr Croatia 10% .eu European Union 2% .ba Bosnia Bulgarian 26% .bg Bulgaria 10% .eu European Union Irish 20% .ie Ireland 10% .eu European Union Finnish 44% .fi Finland 7% .eu European Union
Table 11: Frequent geographic TLDs in the parallel HPLT v2 datasets for different languages (non-English side).
Appendix H Manual quality inspection
Language % Porn $\downarrow$ % Unnat. $\downarrow$ % LID $\uparrow$
Arabic 0 (-) 9 (5-13) 100 (-)
Asturian 0 (-) 28 (22-35) 69 (62-75)
Bengali 1 (-) 0 (-) 100 (-)
Catalan 0 (-) 14 (9-19) 99 (-)
Czech 0 (-) 9 (4-13) 100 (-)
Chinese 0 (-) 25 (18-31) 99 (-)
Dutch 1 (-) 5 (-) 100 (-)
English 1 (-) 13 (8-18) 100 (-)
Finnish 1 (-) 4 (-) 100 (-)
German 1 (-) 2 (-) 98 (-)
Hindi 2 (-) 2 (-) 98 (-)
Iran. Persian 0 (-) 25 (18-31) 99 (-)
Marathi 0 (-) 6 (-) 97 (-)
Modern Greek 0 (-) 3 (-) 100 (-)
Nor. Bokmål 2 (-) 8 (4-11) 99 (-)
Nor. Nynorsk 0 (-) 3 (-) 93 (-)
Polish 1 (-) 7 (3-11) 100 (-)
Russian 2 (1-3) 18 (15-21) 98 (-)
Scot. Gaelic 0 (-) 3 (-) 89 (85-93)
Slovak 0 (-) 10 (6-14) 100 (-)
Spanish 1 (-) 9 (5-13) 100 (-)
Turkish 6 (-) 10 (5-14) 99 (-)
Table 12: Manual quality inspection of a random sample of documents from the cleaned version, stratified by crawls groups. Percentages of extracted texts considered as pornography (Porn), unnatural texts (Unnat.), and texts correctly classified by language identification (LID) (the 95% confidence intervals for the percentage estimates are given in brackets when applicable).
In this section, we study how the quality of the extracted texts varies between older and newer crawls, and also between IA and CC crawls. More specifically, for a particular language we wanted to understand if there are any substantial differences in the proportions of texts classified as this language by mistake or just undesirable texts.
For this study, we carried out manual annotation of documents from the cleaned version of our dataset asking our annotators to provide three binary annotations for each document.
  • LID ok: 0 if most of the text is not in the target language, otherwise 1;
  • Unnatural: 1 if most of the text looks unnatural (e.g. word lists for SEO, mostly boilerplate, etc.), otherwise leave empty;
  • Porn: 1 if the text looks like pornographic content, otherwise leave empty.
We compared four groups of crawls: among wide IA crawls and CC crawls separately we selected old crawls from 2012-2014 and new crawls from 2017-2020. Among languages spoken by the paper authors, 22 languages were selected for annotation.
Figure 11: Proportions of unnatural texts among the cleaned texts extracted from four selected groups of crawls, according to manual inspection of a sample. Error bars correspond to the 95% confidence intervals.
For each language and each group of crawls, 50 random documents from the cleaned version of our datasets were annotated by a native or a fluent speaker of this language. In total, 200 documents for each language were annotated, except for Russian where three native speakers annotated 600 documents. Only texts extracted from the documents were shown to the annotators, they did not know which crawl each text came from or any other meta-information. For documents longer than 1000 characters, the first 500 characters and 500 characters from the beginning of the second half were shown.
Table 12 shows the results for the four groups combined together. [^33] Since the sample is stratified by group and the crawls from these groups give about 52% of all texts in our dataset, one should carefully interpret these statistics in the context of the full dataset. We see that the proportion of pornographic content is low, usually between 0-2% with a maximum of 6% for Turkish. The precision of our LID model for the inspected languages is above 97%, with a few notable exceptions. The worst precision is for Asturian where we observed about 30% of texts being in Spanish or other Spanish minority languages (e.g. Extremeño, Aragonese), or just SEO lists consisting of e.g. song names not in Asturian. The proportions of unnatural texts vary a lot from language to language. Annotators report the following major types of unnaturalness: lists of services and goods, commercial ads with varying degrees of grammaticality, traces of Wikipedia markup, documents consisting mostly of menus, and boilerplate missed by boilerplate removal.
Figure 11 shows proportions of unnatural texts for each language and group of crawls. Looking at individual languages, for most of them the group of new CC crawls give a much lower proportion of unnatural texts than other groups. However, since only 50 documents were labelled from each group and language, the confidence intervals are large and statistically significant conclusions cannot be made for each individual language. However, when annotations for all languages are combined (denoted as TOTAL on the figure) it becomes clear that for a random language (among those annotated) a random document has about 2x lower probability to be unnatural if it comes from the group of newer CC crawls compared to older CC crawls or any of two groups of IA crawls. For the proportions of pornographic content and documents misclassified by LID we did not observe any consistent differences for different groups of crawls.
Appendix I Model training and evaluation
I.1 Corpora comparison: English
Pretraining
We fully replicated the original FineWeb training and evaluation setup by Penedo et al. (2024a), with the same architecture and pretraining settings (1.71B parameters, Llama architecture with a sequence length of 2048 tokens, GPT 2 tokenizer, and a global batch size of ~2 million tokens). We train 4 models that are differentiated only by training data, and evaluate their performance at different stages of model training. Each model is trained on 100 billion tokens, randomly sampled from the following datasets:
  • English HPLT v2 data, cleaned
  • English HPLT v2 data, deduplicated
  • English HPLT v1.2 de Gibert et al. (2024)
  • FineWeb dataset Penedo et al. (2024a)
We use NVIDIA’s Megatron-LM (https://github.com/NVIDIA/Megatron-LM) training framework instead of HuggingFace’s nanotron (https://github.com/huggingface/nanotron) framework used by Penedo et al. (2024a). Each model is trained on the LUMI supercomputer with 16 nodes, each with 4 AMD MI250x GPUs with dual-GCD (graphics compute die) design, amounting to 8 logical devices. In total, we used 128 devices and a single 64-core CPU for approximately 84 hours, totalling 11 008 GPU hours per model.
Evaluation
Evaluation is performed using HuggingFace’s LightEval tool Fourrier et al. (2023) on the tasks listed below. Results per task are presented in Figure 12.
  • HellaSwag: a dataset to evaluate commonsense reasoning. Its questions are designed to be trivial for humans but challenging for LLMs Zellers et al. (2019).
  • PIQA: a dataset focusing on reasoning with multiple-choice questions about physical interactions, evaluating the LLM’s understanding of how different objects are used in various situations Bisk et al. (2020).
  • OpenBookQA: a dataset consisting of multiple-choice questions which require understanding concepts and their relations, benchmarking the complex reasoning and inference performance of the LLM Mihaylov et al. (2018).
  • ARC Easy and ARC Challenge: both parts of the AI2 Reasoning Challenge dataset, containing easier and more complex questions to test the LLM’s reasoning skills Clark et al. (2018).
Figure 12: Final checkpoint scores of the English models trained on the datasets shown, grouped based on the benchmarks conducted. The models perform quite similarly with the exception of the model trained on the HPLT v1.2 dataset, the scores of which are noticeably lower.
I.2 Corpora comparison: Norwegian
Pretraining
We mirrored the pretraining setup used for the English ablation studies in Section I.1, except for two details: 1) we trained a new tokenizer specifically for Norwegian, using a single tokenizer for all experiments trained on equal number of samples from all ablated corpora using the tokenizers library; 2) we pretrained the models for 30B tokens (roughly corresponding to 1 epoch on most of the ablated corpora) instead of 100B, mirroring the multilingual experiments for FineTasks (Kydlíček et al., 2024).
We compared five different filtered corpora that support Norwegian. Most of these discriminate between two written variants of Norwegian – Bokmål and Nynorsk – in those cases, we simply concatenate these subcorpora. The ablated corpora are:
  • Norwegian HPLT v2 data, cleaned;
  • Norwegian CulturaX (Nguyen et al., 2024a);
  • Norwegian HPLT v1.2 (de Gibert et al., 2024);
  • Norwegian FineWeb-2 (Penedo et al., 2024b);
  • Norwegian mC4 Xue et al. (2021).
The pretraining code is built on the Megatron-DeepSpeed framework (Smith et al., 2022). All models were trained on the LUMI supercomputer using 32 compute nodes, each with 4 AMD MI250x GPUs. The full pretraining run of each model took approximately 15 hours (wall-clock time), or 1 920 GPU-hours ($15\times 32\times 4$ hours), respectively.
Evaluation
Evaluation is performed using NorEval Mikhailov et al. (2025a), an open-source benchmark for Norwegian built upon LM Evaluation Harness Gao et al. (2024). We consider the following ten multiple-choice QA, generative QA, sentence completion, and sentence pair ranking tasks that target different aspects of the model understanding and generation abilities in Norwegian Bokmål and Nynorsk:
  • Commonsense reasoning: performing logical and commonsense reasoning (NorCommonsenseQA, Mikhailov et al., 2025b).
  • Norwegian-specific & world knowledge: answering questions about facts and Norwegian culture (NorOpenBookQA and NRK-Quiz-QA, Mikhailov et al., 2025b).
  • Norwegian language knowledge: understanding Norwegian punctuation rules (NCB, Mikhailov et al., 2025a) [^34] https://huggingface.co/datasets/hcfa/ncb and idioms (NorIdiom, Mikhailov et al., 2025a). [^35] https://huggingface.co/datasets/Sprakbanken/Norwegian_idioms
  • Machine reading comprehension: understanding a given text and extracting an answer from it (NorQuAD, Ivanova et al., 2023).
We aim to find tasks that provide a reliable signal during pretraining. We evaluate the models in a zero-shot regime at regular checkpoint intervals (approx. 1B tokens) on all tasks. Next, we discard tasks that provide a low signal based on two criteria Penedo et al. (2024b):
  • Monotonicity: the Spearman rank correlation between the number of steps and the target performance score is at least 0.5 across all model checkpoints.
  • Non-random performance: the difference between the random baseline (zero for generative tasks, one divided by the number of answer choices for multiple-choice tasks, and a coin flip probability for sentence pair ranking tasks) and the maximum score across all models is positive and satisfactory.
Figure 13: Final checkpoint scores of the Norwegian models trained on the datasets shown, grouped based on the evaluation datasets in NorEval. The models perform quite similarly with the exception of the model trained on the HPLT v1.2 dataset, the NorQuAD and NCB scores of which are generally lower.
The filtering results in four datasets: NCB (accuracy), NRK-Quiz-QA Bokmål (accuracy), NorCommonsenseQA Bokmål (accuracy), and NorQuAD (F1-score). We aggregate the performance across the datasets using the average normalized score Myrzakhan et al. (2024). We report the performance results for our 150 checkpoints in Figure 5 (see §6.2) and final checkpoint performance in Figure 13.
Appendix J LTG-BERTs training and evaluation details
Following the HPLT v1.2, [^36] https://hplt-project.org/HPLT_D4_1___First_language_models_trained.pdf we use UD treebanks of version 2.13373737https://lindat.mff.cuni.cz/repository/xmlui/handle/11234/1-5287 for most languages, except for Albanian and Georgian. These languages were not used in the HPLT v1.2 report due to missing training and development splits in UD 2.13. However, UD 2.15 does contain the required splits, and we use them. We do not evaluate NER on Maltese, since its WikiAnn training split contains only 100 samples. Table 13 shows detailed MLM evaluation results by language and task.
LTG-BERT architecture Samuel et al. (2023) is a version of the original masked BERT model Devlin et al. (2019). The differences include removing the next sentence prediction objective, swapping subword masking to span masking, and other minor architectural improvements. LTG-BERT was shown to perform well for small-sized training datasets Samuel (2023), which fits our evaluation setup. The models were trained with the same hyperparameters as in the aforementioned HPLT report.
We trained separate models for Bosnian and Croatian, in addition to the joint Bosnian-Croatian model. Since the UD does not provide Bosnian treebanks, we evaluated all three models on the Croatian datasets. We did not include Serbian, because it uses the Cyrillic writing system in HPLT v2, while UD features Serbian data only in Latin. Exploring whether mixing the scripts still improves the results is left for future work. It is difficult to give any clear recommendations on which of the three models to use for practical tasks, since all of them yield satisfactory evaluation results (ranking varies from task to task).
LTG-BERT models were trained for 31 250 steps on 16 compute nodes with 4 physical AMD INSTINCT MI250x GPUs each for approximately 9.8 hours. Sharding, training a tokenizer and tokenizing for larger languages required up to 3.5, 0.5 and 1 hours correspondingly on 7 AMD EPYC 7763 CPUs (these numbers are estimated from the processing of English, the largest data subset in HPLT v2. The processing time of different languages may vary, for instance, languages without whitespace separation between words require an additional pretokenizing step). UD fine-tuning and NER fine-tuning required 1.1 hours and 8 minutes correspondingly on 1 GPU (estimated for English).
POS tags Lemmas Dependency parsing NER Language mBERT XLM-R HPLT HPLT v2 mBERT XLM-R HPLT HPLT v2 mBERT XLM-R HPLT HPLT v2 mBERT XLM-R HPLT HPLT v2 v1.2 v1.2 v1.2 v1.2 als_Latn 59.1 61.6 64.0 64.5 78.2 75.0 76.3 77.2 33.1 29.3 25.3 24.7 92.3 92.9 92.4 93.9 bel_Cyrl 94.1 94.6 95.5 95.7 93.2 93.8 93.8 97.1 88.1 89.9 91.1 91.7 91.7 90.3 90.1 92.8 bos_Latn 95.5 96.2 96.4 96.6 97.2 97.4 97.2 97.1 90.2 91.3 91.3 91.7 91.5 91.6 89.3 92.8 hrv_Latn 95.5 96.2 96.4 96.8 97.2 97.4 97.2 97.1 90.2 91.3 91.3 91.6 91.5 91.6 89.3 92.5 bul_Cyrl 97.0 97.5 97.8 97.9 97.5 97.7 97.3 97.3 92.7 94.4 94.0 94.5 92.2 92.2 91.5 93.0 cat_Latn 97.1 97.2 97.4 97.5 99.4 99.4 99.4 97.5 93.6 94.1 94.4 99.4 92.1 91.0 90.1 94.5 ces_Latn 97.8 98.0 98.3 98.4 99.3 99.3 99.4 99.4 93.5 94.2 94.4 94.6 91.2 91.2 89.0 91.8 cym_Latn 87.2 88.3 89.2 89.0 94.6 94.4 93.7 92.3 80.8 82.8 82.3 82.8 92.5 90.0 89.4 93.4 dan_Latn 96.7 97.8 97.8 97.9 97.2 97.6 97.1 97.1 86.7 89.1 88.8 89.5 91.2 91.6 90.3 92.0 deu_Latn 88.8 89.4 80.7 89.9 97.6 97.7 95.5 97.5 84.6 87.1 76.4 87.6 89.4 87.7 64.1 89.2 ell_Grek 94.6 95.7 96.1 96.2 94.6 94.7 94.1 94.1 91.7 93.5 92.2 93.2 90.2 90.7 90.2 92.6 eng_Latn 96.1 96.8 96.7 97.0 97.8 98.0 97.9 98.1 91.3 92.6 92.2 93.0 2.2 81.1 81.0 82.7 spa_Latn 95.7 95.9 96.0 96.2 99.4 99.4 99.4 99.4 92.3 93.0 93.1 93.4 90.9 89.9 89.6 90.8 est_Latn 96.0 96.6 97.1 97.1 94.8 95.0 95.2 95.2 88.1 89.7 90.8 91.0 91.8 90.4 89.6 93.0 eus_Latn 91.0 91.4 92.3 92.3 95.7 95.9 96.0 95.9 85.3 87.3 88.1 88.2 91.3 90.7 89.8 92.9 pes_Arab 95.9 96.3 96.4 96.3 99.1 99.4 99.4 99.5 92.7 93.8 93.9 94.1 92.0 92.9 91.8 93.9 fin_Latn 95.1 96.4 96.8 97.0 90.6 91.5 91.6 91.4 90.2 93.0 93.3 94.0 90.2 90.0 89.2 91.6 fra_Latn 97.8 98.1 98.1 98.0 98.6 98.8 93.8 98.6 93.8 94.4 94.5 94.8 90.5 88.7 87.2 90.0 gle_Latn 86.5 87.1 88.7 89.3 95.5 95.8 96.1 95.6 81.3 82.7 83.4 84.3 80.8 78.0 55.9 78.2 glg_Latn 96.9 97.1 97.1 97.0 98.3 98.3 98.2 98.0 82.3 82.6 82.3 82.2 92.5 93.3 91.1 94.1 heb_Hebr 95.6 96.1 96.5 96.7 97.0 97.2 97.1 97.2 89.8 91.6 91.0 91.9 2.6 84.2 88.4 89.3 hin_Deva 92.4 93.3 93.6 93.7 98.9 99.0 99.0 99.0 92.6 93.3 93.5 93.6 88.6 88.0 84.3 89.5 hrv_Latn 95.5 96.2 96.4 96.7 97.2 97.4 97.2 97.2 90.2 91.3 91.3 91.8 91.5 91.6 89.3 92.0 hun_Latn 93.0 94.3 93.0 94.1 93.0 94.3 93.0 92.3 84.3 86.7 82.4 86.1 92.2 91.9 92.8 93.1 hye_Armn 88.7 91.2 92.7 92.7 94.4 94.9 93.9 94.7 80.4 85.3 84.1 86.8 95.7 95.3 94.8 95.9 ind_Latn 89.5 89.8 89.6 89.1 98.2 98.3 98.0 97.5 82.4 82.7 81.7 81.8 91.3 91.6 89.1 92.0 isl_Latn 87.7 88.1 88.6 88.7 96.2 96.4 96.5 96.4 85.2 86.6 86.9 87.4 81.7 63.9 55.9 78.3 ita_Latn 98.0 98.0 98.1 98.3 98.6 98.7 98.8 98.7 94.1 94.4 94.6 95.1 90.5 89.7 87.8 91.2 jpn_Jpan 97.5 97.7 97.8 97.8 98.3 98.3 98.3 98.4 94.1 94.6 94.6 94.8 66.5 65.9 67.4 67.2 kat_Geor 91.3 92.6 92.4 92.4 92.8 93.7 92.5 92.5 79.5 80.9 80.8 81.3 87.2 4.7 89.6 90.7 kor_Hang 88.6 89.7 89.9 90.1 94.0 94.3 94.4 94.4 88.0 89.0 89.4 89.7 87.8 87.0 88.3 89.3 lvs_Latn 91.6 92.8 92.4 93.6 96.9 91.6 96.8 97.7 88.8 90.9 90.9 92.1 93.2 92.6 90.7 93.9 lit_Latn 87.7 91.9 92.0 92.5 90.2 91.6 91.5 91.2 79.3 85.7 84.9 86.8 89.1 89.3 87.0 91.0 ltz_Latn - - - - - - - - - - - - - - - 89.2 mkd_Cyrl - - - - - - - - - - - - - - - 94.6 mlt_Latn 94.7 94.5 97.0 97.7 100.0 100.0 100.0 100.0 78.2 78.5 83.2 87.2 - - - - nob_Latn 97.0 97.4 97.6 97.5 98.5 98.8 98.8 98.7 93.2 94.3 94.5 94.7 91.9 92.6 91.1 93.2 nld_Latn 96.2 96.9 97.1 97.2 94.1 94.7 94.4 94.1 91.6 92.9 93.8 94.1 91.7 90.4 88.6 91.0 nno_Latn 96.6 97.0 97.7 97.8 98.2 98.4 98.5 98.5 92.9 93.9 94.6 95.0 95.8 93.6 93.2 95.5 pol_Latn 95.6 95.5 96.9 97.2 97.8 98.2 98.2 98.2 93.7 95.2 95.3 95.6 12.9 88.8 89.7 89.6 por_Latn 93.6 94.0 94.1 94.1 98.1 98.3 98.3 98.2 83.4 84.5 84.9 85.3 91.2 90.3 88.0 91.5 ron_Latn 97.3 97.6 97.7 97.9 97.7 97.9 97.8 97.8 89.5 91.0 90.6 91.6 94.5 93.6 91.2 93.6 rus_Cyrl 93.8 94.4 94.5 94.7 98.3 98.5 98.6 98.6 92.6 93.4 93.6 93.8 88.0 86.9 85.6 89.0 slk_Latn 89.1 97.6 98.1 91.9 95.7 96.1 95.6 95.5 92.9 94.4 93.8 95.0 93.2 92.9 91.2 93.3 slv_Latn 96.7 97.6 98.1 98.2 98.5 98.7 98.6 98.7 93.4 94.7 94.8 95.3 93.4 93.1 93.6 94.2 srp_Cyrl - - - - - - - - - - - - 91.6 92.4 - 93.4 swe_Latn 96.5 97.4 97.4 97.3 97.3 97.6 97.1 97.0 89.4 92.1 90.8 91.7 94.3 94.5 93.5 94.4 tat_Cyrl - - - - - - - - - - - - 89.7 80.6 82.9 84.0 tur_Latn 90.4 91.0 91.5 91.4 91.1 91.3 91.9 91.4 70.9 73.0 73.6 74.6 92.2 92.0 90.8 92.5 ukr_Cyrl 93.1 94.7 72.9 95.3 87.0 97.2 87.0 97.0 89.4 91.8 61.3 92.1 92.0 91.7 77.5 92.8 vie_Latn 89.8 92.1 91.8 92.1 99.9 99.9 99.9 99.9 66.5 70.3 68.0 70.3 91.9 90.6 89.2 90.3 zho_Hans 96.2 96.3 96.0 96.0 99.9 99.9 99.9 99.9 86.1 86.9 84.6 85.6 0.1 76.5 75.5 74.5
Table 13: Results of monolingual masked language models trained on the HPLT v2 datasets compared to the baselines on part-of-speech (POS) tagging, lemmatization, dependency parsing and named entity recognition. For POS tagging, we evaluate the AllTags performance, which is the exact match accuracy of the UPOS, XPOS and UFeats UDtags. For dependency parsing, we report LAS, and for lemmatization accuracy.
Appendix K Full Results for Translation Models Built on Parallel Data
We compare models trained on HPLT v2, Tatoeba (Tiedemann, 2012, 2020), and the combination of the two datasets. The language selection is the intersection of the languages covered by both datasets. We evaluate the models on the FLORES-200 evaluation benchmark (NLLB Team et al., 2024) using SacreBLEU implementation of BLEU383838nrefs:1|case:mixed|eff:no|smooth:exp|version:2.5.1,
and where applicable, tok:ja-mecab, tok:ko-mecab, or tok:13a
and chrF++393939nrefs:1|case:mixed|eff:yes|nc:6|nw:0|space:no|version:2.5.1 metrics (Post, 2018) and COMET-22-DA (Rei et al., 2022).
Tables 14 and 15 present the full results of the MT models for translation into English and from English respectively. For reference, we also include the performance of models trained on the HPLT v1.2 dataset, which shares the same underlying extraction pipeline. Note that we did not perform any language-specific hyper-parameter tuning which possibly led to low scores for a few model instances.
HPLT v1.2 HPLT v2 Tatoeba (OPUS) HPLT v2+OPUS BLEU chrF++ COMET BLEU chrF++ COMET BLEU chrF++ COMET BLEU chrF++ COMET en-af 39.2 64.5 0.8398 38.3 63.6 0.8397 38.8 63.8 0.8409 en-ar 26.6 55.0 0.8442 en-az 12.2 41.0 0.8128 11.3 38.7 0.8074 11.5 38.7 0.8011 en-be 11.6 39.0 0.7756 11.2 37.4 0.7767 11.8 38.2 0.7800 en-bg 38.0 62.4 0.8680 0.9 14.5 0.6774 30.0 51.9 0.8122 en-bn 16.0 45.2 0.8109 16.6 45.9 0.8282 16.8 46.1 0.8275 en-bs 4.7 26.0 0.4314 26.8 53.9 0.8672 en-ca 38.4 61.7 0.8461 37.8 61.0 0.8334 39.8 62.2 0.8440 39.5 62.1 0.8425 en-cy 50.4 69.9 0.8611 47.7 67.6 0.8536 48.4 67.8 0.8506 en-eo 27.2 54.7 0.8264 25.5 54.1 0.8500 25.9 54.4 0.8523 en-et 23.7 53.4 0.8664 24.5 53.8 0.8684 24.5 53.3 0.8600 24.4 53.3 0.8578 en-eu 12.1 43.4 0.7674 16.5 49.5 0.8215 14.9 47.2 0.8098 14.8 47.1 0.8122 en-fa 21.5 47.5 0.7947 23.4 50.0 0.8336 23.6 50.0 0.8349 en-fi 21.3 51.1 0.8709 21.3 51.2 0.8725 22.1 51.6 0.8752 en-ga 27.3 52.6 0.7561 29.0 53.9 0.7543 30.2 53.9 0.7715 30.8 54.6 0.7717 en-gl 27.9 54.0 0.8033 30.0 55.7 0.8179 31.4 56.1 0.8302 31.4 56.1 0.8264 en-gu 19.3 46.5 0.8066 22.5 49.9 0.8518 22.6 49.9 0.8479 en-he 28.1 54.0 0.8320 29.7 55.9 0.8532 29.6 55.4 0.8503 en-hi 32.8 55.5 0.7621 32.0 54.6 0.7612 33.1 55.5 0.7728 32.5 54.9 0.7658 en-hr 27.5 53.7 0.8504 en-is 20.6 45.1 0.7651 22.2 47.1 0.7766 22.8 47.1 0.7800 23.1 47.5 0.7859 en-ja 27.0 26.6 0.8244 29.9 30.2 0.8640 29.6 29.9 0.8633 en-kk 21.0 51.4 0.8651 16.5 45.1 0.8315 16.9 45.3 0.8347 en-kn 13.8 43.5 0.7746 19.5 50.8 0.8348 19.2 51.1 0.8369 en-ko 25.0 31.2 0.8268 26.6 32.2 0.8424 26.5 32.0 0.8402 en-lt 22.9 50.9 0.8214 25.0 52.7 0.8486 24.4 52.5 0.8417 en-lv 26.8 53.1 0.8214 23.9 50.0 0.7898 24.3 50.6 0.7891 en-mk 32.2 58.6 0.8478 33.9 60.1 0.8718 34.1 60.1 0.8660 en-ml 0.6 20.2 0.5753 14.4 47.9 0.8438 14.6 48.0 0.8427 en-mr 11.0 37.9 0.6086 13.9 42.3 0.6808 14.2 42.4 0.6792 en-ms 38.3 63.9 0.8580 24.4 52.6 0.8534 25.1 53.2 0.8540 en-mt 34.7 63.2 0.7072 35.8 64.1 0.7102 36.0 64.4 0.7111 en-nb 33.0 58.2 0.8661 en-ne 0.9 20.9 0.4880 12.9 42.8 0.7404 12.8 43.1 0.7386 en-nn 22.8 47.1 0.7766 en-si 1.2 18.6 0.6289 13.1 41.0 0.8542 13.2 41.2 0.8548 en-sk 29.3 54.0 0.8279 29.3 53.9 0.8353 29.9 54.5 0.8423 en-sl 26.8 52.2 0.8295 26.5 52.0 0.8339 27.5 52.6 0.8414 en-sq 27.8 54.6 0.8509 27.7 54.2 0.8398 29.9 55.8 0.8659 29.3 55.3 0.8600 en-sr 32.2 57.5 0.8512 en-sw 28.4 54.6 0.7743 32.5 58.2 0.7964 31.2 57.0 0.8058 31.3 57.0 0.8031 en-ta 14.1 46.7 0.8139 16.7 50.3 0.8565 16.4 50.1 0.8553 en-te 20.2 51.3 0.8104 22.1 53.7 0.8378 22.7 53.9 0.8383 en-th 9.9 40.9 0.7977 8.1 40.6 0.8053 8.7 40.9 0.8053 en-tr 25.3 53.7 0.8368 27.8 56.4 0.8685 27.5 55.8 0.8638 en-uk 26.7 52.6 0.8457 27.2 53.4 0.8532 26.8 52.8 0.8471 en-ur 18.9 43.2 0.7548 19.3 44.0 0.7537 19.5 44.6 0.7584 en-uz 16.3 49.1 0.8397 15.9 47.3 0.8497 17.1 48.8 0.8532 en-vi 37.8 55.8 0.8358 39.3 57.1 0.8489 38.8 56.6 0.8451 en-xh 12.0 44.2 0.7323 0.0 3.5 0.2317 0.0 3.5 0.1647
Table 14: MT results (BLEU, chrf++, COMET22-DA) for models translating from English, trained on our HPLT v2, Tatoeba (OPUS), a combination of both, and the existing HPLT v1.2 (numbers reported where available).
HPLT v1.2 HPLT v2 Tatoeba (OPUS) HPLT v2+OPUS BLEU chrF++ COMET BLEU chrF++ COMET BLEU chrF++ COMET BLEU chrF++ COMET az-en 18.5 47.1 0.8290 17.4 44.7 0.8039 18.6 46.2 0.8175 be-en 16.1 46.7 0.7886 14.8 44.7 0.7621 15.5 45.8 0.7743 bg-en 35.5 61.1 0.8556 7.4 32.5 0.5104 34.8 60.6 0.8524 bn-en 27.9 53.7 0.8468 28.4 53.7 0.8498 29.0 54.2 0.8523 bs-en 12.8 38.0 0.5882 35.2 59.9 0.8508 ca-en 41.0 64.4 0.8676 39.2 63.1 0.8478 41.1 64.4 0.8580 40.3 63.9 0.8541 cy-en 51.5 70.7 0.8615 50.0 68.8 0.8456 50.6 69.1 0.8455 eo-en 35.5 59.2 0.8362 35.6 59.2 0.8437 35.6 59.3 0.8429 et-en 30.6 56.6 0.8611 30.3 55.8 0.8517 30.7 56.1 0.8510 30.7 55.6 0.8510 eu-en 19.4 45.7 0.7810 23.3 49.2 0.8121 22.2 47.5 0.8064 22.0 47.4 0.8042 fa-en 31.1 56.4 0.8447 33.7 58.2 0.8585 32.7 57.6 0.8546 fi-en 26.6 52.4 0.8442 26.6 51.9 0.8451 26.2 51.6 0.8421 ga-en 29.9 54.9 0.7653 34.1 58.8 0.8006 32.3 56.3 0.7754 33.1 57.7 0.7918 gl-en 31.4 57.2 0.8236 33.7 59.2 0.8374 34.5 59.1 0.8395 35.0 59.9 0.8441 gu-en 28.5 54.6 0.8475 32.0 57.0 0.8646 33.0 57.6 0.8667 he-en 38.2 62.2 0.8534 39.7 62.9 0.8622 40.4 63.6 0.8665 hi-en 35.2 59.9 0.8741 34.7 59.5 0.8701 35.8 60.1 0.8738 36.9 61.0 0.8773 hr-en 31.7 56.4 0.8389 is-en 25.3 50.0 0.7815 29.0 53.4 0.8189 29.0 52.8 0.8163 28.7 52.8 0.8136 ja-en 19.9 46.8 0.8255 24.6 52.5 0.8628 23.6 50.6 0.8533 kk-en 27.0 53.4 0.8403 22.6 47.8 0.8003 22.6 47.7 0.7998 kn-en 3.8 24.5 0.6246 27.9 53.4 0.8391 27.4 53.2 0.8396 ko-en 24.1 51.3 0.8458 25.7 52.7 0.8586 25.8 52.7 0.8578 lt-en 26.5 51.9 0.8138 27.2 52.6 0.8224 27.0 52.2 0.8201 lv-en 29.3 56.0 0.8368 25.0 50.9 0.7862 26.6 53.3 0.8113 mk-en 37.0 61.4 0.8522 38.2 62.0 0.8558 38.8 62.6 0.8596 ml-en 2.9 23.5 0.5978 26.4 51.7 0.8342 26.4 51.9 0.8363 mr-en 23.8 49.8 0.8063 26.1 51.9 0.8299 26.9 52.2 0.8320 ms-en 37.2 61.3 0.8561 38.5 61.8 0.8579 38.0 61.7 0.8583 mt-en 45.0 68.4 0.7777 47.1 68.5 0.7892 47.5 68.9 0.7895 nb-en 37.3 61.0 0.8540 ne-en 22.1 47.9 0.7929 26.6 52.0 0.8408 27.7 52.8 0.8436 nn-en 32.9 55.8 0.7945 si-en 3.0 24.2 0.5979 26.0 51.2 0.8381 26.9 51.9 0.8418 sk-en 31.6 58.1 0.8456 32.7 58.6 0.8486 33.2 59.0 0.8535 sl-en 29.2 55.0 0.8371 28.7 54.4 0.8345 29.7 55.6 0.8402 sq-en 31.7 58.3 0.8468 32.1 58.6 0.8453 33.7 58.8 0.8448 34.8 59.8 0.8534 sr-en 37.4 62.7 0.8544 sw-en 27.2 51.0 0.7542 35.3 57.8 0.8086 34.3 56.3 0.7979 33.5 55.6 0.7932 ta-en 24.0 49.6 0.8068 23.4 49.0 0.8139 24.3 49.5 0.8154 te-en 30.2 55.3 0.8328 31.5 55.9 0.8438 31.9 56.4 0.8446 th-en 24.9 52.3 0.8452 22.9 51.0 0.8381 23.7 51.7 0.8410 tr-en 29.5 54.9 0.8392 32.2 57.3 0.8622 32.7 57.4 0.8602 uk-en 33.1 58.7 0.8444 33.4 59.2 0.8470 33.9 59.6 0.8478 ur-en 26.3 52.1 0.8138 26.2 50.9 0.8097 27.4 52.0 0.8144 uz-en 24.8 51.5 0.8110 23.6 48.6 0.7990 24.8 50.0 0.8064 vi-en 32.0 56.4 0.8514 33.5 57.9 0.8602 32.9 57.2 0.8543 xh-en 22.1 45.5 0.6703 0.1 11.8 0.2871 0.2 13.1 0.2628
Table 15: MT results (BLEU, chrf++, COMET22-DA) for models translating into English, trained on our HPLT v2, Tatoeba (OPUS), a combination of both, and the existing HPLT v1.2 (numbers reported where available).